### v1

# Ehrhart quasi-polynomials via Barnes polynomials and discrete moments of parallelepipeds

Abstract
We give novel and explicit formulas for the Ehrhart quasi-polynomials of rational simple polytopes, in terms of Barnes polynomials and discrete moments of half-open parallelepipeds. These formulas also hold for all positive dilations of a rational polytope. There is an interesting appearance of an extra complex $z$-parameter, which seems to allow for more compact formulations.
We also give similar formulas for discrete moments of rational polytopes, and their positive dilates, objects known in the literature as sums of polynomials over a polytope. The appearance of the Barnes polynomials and the Barnes numbers allow for explicit computations. From this work, it is clear that the complexity of computing Ehrhart quasi-polynomials lies mainly in the computation of various discrete moments of parallelepipeds. These discrete moments are in general summed over a particular lattice flow on a closed torus, defined in this paper.
Some of the consequences involve novel vanishing identities for rational polytopes, novel formulations of Ehrhart polynomials of unimodular polytopes, and a differential equation that extends the work of Eva Linke.
1 Introduction, and statements of the main results
One of the most natural definitions for a discrete volume of a rational polytope ${\mathcal{P}}\subset\mathbb{R}^{d}$ is its classical integer point enumerator $L_{\mathcal{P}}(t):=|\mathbb{Z}^{d}\cap t{\mathcal{P}}|$, where $t{\mathcal{P}}$ is the $t$-dilate of $\mathcal{P}$}. When $\mathcal{P}$} is an integer (resp. rational) polytope, and $t$ is a positive integer, it is known by Ehrhart’s original work [15] that $L_{\mathcal{P}}(t)$ is a polynomial (resp. quasi-polynomial) in the integer-valued parameter $t$.
Here we give novel formulas for Ehrhart quasi-polynomials of rational simple polytopes, in terms of Barnes polynomials and discrete moments of half-open parallelepipeds. These formulas hold for all positive values of $t$, and for any rational polytope.
We find here two main objects that form the complete building blocks for the Ehrhart quasi-polynomials of rational polytopes: the collection of Barnes polynomials $B_{k}(t,{\mathbf{a}})$, and the discrete moments of fundamental parallelepipeds, defined in (12) below.
The Barnes polynomials are defined by their generating function:
$\frac{x^{d}e^{tx}}{(e^{a_{1}x}-1)\cdots(e^{a_{d}x}-1)}:=\sum_{k\geq 0}B_{k}(t,{\mathbf{a}})\frac{x^{k}}{k!},$ (1)
where initially ${\mathbf{a}}:=(a_{1},\dots,a_{d})\in\mathbb{C}^{d}$ is any fixed vector of complex numbers whose real part is positive, where $t\in\mathbb{\mathbb{R}}_{>0}$, and where the identity (1) is valid for all $x\in\mathbb{C}$ inside the domain of convergence of the series. It follows easily from (1) that for each nonnegative integer $k$, $B_{k}(t,{\mathbf{a}})$ is a polynomial function of $t$. It also follows that when $d=1$, $B_{k}(t,(1))$ is the classical Bernoulli polynomial. Because the Barnes polynomials are a natural multivariable extension of the classical Bernoulli polynomials, these two families of polynomials enjoy some similar properties. We note that some papers refer to the polynomials $B_{k}(t,{\mathbf{a}})$ as Bernoulli-Barnes polynomials, while others refer to them as multiple Bernoulli polynomials. In $1901$ E. W. Barnes [3] initiated the systematic study of the polynomials $B_{k}(t,{\mathbf{a}})$. Barnes also related these polynomials to multiple gamma functions, and to some useful zeta functions. In recent years, the Barnes polynomials have been studied extensively by the important work of Bayad et al. [6, 7, 8].
We use elementary techniques, and our first main result is an explicit formula for the Ehrhart polynomial (resp. quasi-polynomial) of any simple integer (resp. rational) polytope $\mathcal{P}$}, in terms of Barnes polynomials and discrete moments of parallelepipeds. In this sense, this paper also initiates the study of discrete moments of parallelepipeds and more generally the discrete moments of rational polytopes. Such discrete moments may be thought of as belonging to spatial statistics over the integer lattice [13].
For our applications here, we note that if all of the $a_{k}$’s are rational numbers, then definition (1) gives us a polynomial $B_{k}(t,{\mathbf{a}})$ function of $t$, with rational coefficients (see (33) below).
Throughout, we write $V$ for the set of vertices of a polytope $\mathcal{P}$}, and we define the usual local information at each vertex $v$ of a simple rational polytope $\mathcal{P}$}. We call an integer vector $w\in\mathbb{R}^{d}$ primitive if the gcdgcd of all of its coordinates equals $1$. We let
$w_{1}(v),\dots,w_{d}(v)$
be the (unique) primitive integer edge vectors that emanate from a vertex $v\in V$, and let $M_{v}$ be the matrix whose rows are the $w_{k}(v)$’s. At each vertex $v$ of $\mathcal{P}$}, we define the half-open fundamental parallelepiped, with a vertex at the origin, by:
$\Pi_{v}+v:=\{\sum_{k=1}^{d}\lambda_{k}w_{k}(v)\mid 0\leq\lambda<1\}+v,$ (2)
whose volume is equal to $\operatorname{vol}\Pi_{v}=|\det M_{v}|$. We note that $\Pi_{v}$ is a half-open integer parallelepiped with one vertex at the origin.
For any $z\in\mathbb{C}^{d}$, and $x\in\mathbb{R}^{d}$, we will use the slightly unusual, complexified inner product $\langle x,z\rangle:=\sum_{m=1}^{d}x_{m}z_{m}$, noting the absence of complex conjugates in this definition. Given a vertex $v\in{\mathcal{P}}$, we define the vector of complex linear forms
${\mathbf{a}}_{v}:=M_{v}z=\begin{pmatrix}\langle w_{1}(v),z\rangle\\ \vdots\\ \langle w_{d}(v),z\rangle\end{pmatrix}\in\mathbb{C}^{d},$ (3)
which is a vector whose $k$’th coordinate is the inner product $\langle w_{k},z\rangle$.
The starting point for some of our proofs is Brion’s fundamental result for the integer point transform of a rational polytope $\mathcal{P}$}, defined by
$\sigma_{{\mathcal{P}}}(z):=\sum_{q\in{\mathcal{P}}\cap\mathbb{Z}^{d}}e^{\langle q,z\rangle}.$ (4)
The integer point transform $\sigma_{{\mathcal{P}}}(z)$ may be thought of as a discrete version of the Fourier transform [37], and it is a polynomial in the variables $e^{z_{1}},\dots,e^{z_{d}}$, encoding precisely all the integer points in ${\mathcal{P}}\cap\mathbb{Z}^{d}$. In [36], we showed that for integer polytopes $\mathcal{P}$}, the integer point transform is a complete invariant of $\mathcal{P}$}. The discrete Brion Theorem [11] reduces the integer point transform of a rational polytope $\mathcal{P}$} to a sum of the integer point transforms of the vertex tangent cones of $\mathcal{P}$}:
$\sigma_{{\mathcal{P}}}(z)=\sigma_{\mathcal{K}_{v_{1}}}(z)+\cdots+\sigma_{\mathcal{K}_{v_{N}}}(z),$ (5)
where the vertices of $\mathcal{P}$} are $v_{1},\dots v_{N}$ [9]. It turns out, in an elementary way, that each of the integer point transforms $\sigma_{\mathcal{K}_{v}}(z)$ is a rational function of the variables $s_{1}:=e^{z_{1}},\dots s_{d}:=e^{z_{d}}$:
$\sigma_{\mathcal{K}_{v}(z)}:=\sum_{n\in\mathcal{K}_{v}\cap\mathbb{Z}^{d}}e^{\langle n,z\rangle}=\frac{\sigma_{\Pi_{v}+v}(z)}{\prod_{k=1}^{d}\left(1-e^{\langle z,\omega_{k}\rangle}\right)},$ (6)
for each vertex $v\in{\mathcal{P}}$. As a consequence, Brion’s identity (5) is valid for almost all $z\in\mathbb{C}$. More precisely, (5) holds for all $z$ in the set
$\left\{z\in\mathbb{C}^{d}\mid\langle z,w(v)\rangle\notin 2\pi i\mathbb{Z},\text{ for any edge }w(v)\text{ of }{\mathcal{P}}\right\}.$ (7)
Throughout, we will state our theorems with the words “for almost all $z\in\mathbb{C}^{d}$” to mean that they hold for all $z$ in the set (7) above. We also recall that a quasi-polynomial is a function $f(t)=\sum_{k=0}^{d}c_{k}(t)t^{k}$, where each coefficient $c_{k}(t)$ is a periodic function of $t\in\mathbb{Z}_{>0}$. Throughout, even after we extend the domain of $c_{k}(t)$ and hence $f(t)$ to all positive real dilations $t\in\mathbb{R}_{>0}$, as in Linke’s paper [28], we will continue to call $f(t)$ a quasi-polynomial.
It is extremely useful to define the Barnes numbers $B_{k}({\mathbf{a}})$ to be the constant term of the Barnes polynomials $B_{k}(t,{\mathbf{a}})$:
$B_{k}({\mathbf{a}}):=B_{k}(0,{\mathbf{a}}),$ (8)
valid for any ${\mathbf{a}}\in\mathbb{C}^{d}$. This definition is the analogue of the classical $1$-variable Bernoulli numbers. In fact a useful identity for the Barnes numbers in terms of the classical Bernoulli numbers is
$B_{k}({\mathbf{a}})=\sum_{m_{1}+\cdots+m_{d}=k}\binom{k}{m_{1},\dots,m_{d}}a_{1}^{m_{1}-1}\cdots a_{d}^{m_{d}-1}B_{m_{1}}\cdots B_{m_{d}},$ (9)
where the $B_{j}$’s are the classical Bernoulli numbers [7, Theorem 2]. In section 2 below, we give some additional background and examples of Barnes polynomials and Barnes numbers.
The Barnes numbers extend the usual Bernoulli numbers, in the precise sense that for dimension $d=1$, and for ${\mathbf{a}}=(1)$, we have $B_{j}({\mathbf{a}})=B_{j}$, the $j$’th Bernoulli number. A very useful explicit polynomial representation for the Barnes polynomials, in terms of the Barnes numbers $B_{k}({\mathbf{a}})$ is:
$B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j},$ (10)
for each $k\geq 0$.
There is natural flat torus, defined by $\mathbb{R}^{d}/\Lambda$, where $\Lambda$ is the lattice generated by the vertices of any half-open parallelepiped $\Pi$. We’ll identify this torus with the half-open parallelepiped $\Pi$. We define the integer lattice flow over the torus $\mathbb{R}^{d}/\Lambda$, for a the fixed direction $v\in\mathbb{R}^{d}$, by
$LatticeFlow(t):=\mathbb{Z}^{d}-tv\pmod{\Pi}$, (11)
as in Figure 2, Section 4 below. We will compute in a straightforward way some explicit lattice flows over half-open parallelepipeds (see Sections 3 and 4).
Definition 1.
We define the $k$’th discrete moment of a half-open integer parallelepiped $\Pi\subset\mathbb{R}^{d}$ by
$\mu_{k}(\Pi,t):=\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi)}\langle p,z\rangle^{k},$ (12)
for each fixed integer $k\geq 0$, $v\in\mathbb{R}^{d}$ and $t\geq 0$.
Theorem 1.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional, simple rational polytope.
  1. (a)
    The quasi-polynomial $L_{\mathcal{P}}(t):=|t{\mathcal{P}}\cap Z^{d}|$ is equal to:
    $L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k},$ (13)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
  2. (b)
    The coefficient of $t^{r}$ in $L_{\mathcal{P}}(t):=\sum_{r=0}^{d}c_{r}(t)t^{r}$ is equal to:
    $c_{r}(t)=\frac{(-1)^{d}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-r-j},$ (14)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$. The Barnes polynomials $B_{d}(x,{\mathbf{a}})$ are defined in (1), the Barnes numbers $B_{j}({\mathbf{a}})$ are defined in (8), and the vector of linear forms ${\mathbf{a}}_{v}$ is defined in (3). (Proof) $\square$
Remark 1.
It may seem rather surprising a priori that the right-hand-sides of both (16) and (17) appears to depend on $z\in\mathbb{C}^{d}$; but in fact a posteriori they are independent of $z$, as the left-hand-side of (16) shows. Moreover, it is possible that judicious choices of $z$ may reduce the computational complexity of the ensuing discrete moments, and therefore of the Ehrhart polynomials and quasi-polynomials.
Remark 2.
We observe that for rational simple polytopes, one novelty in Theorem 1 is that it reduces the study of Ehrhart quasi-polynomial $L_{\mathcal{P}}(t):=|t{\mathcal{P}}\cap\mathbb{Z}^{d}|$ to the study of discrete moments of integer points in the rational parallelepipeds $\Pi_{v}$, and to the lattice flows (defined in (11)) over its fundamental parallelepipeds.
Moreover, in the case of rational polytopes, the periodicity of the quasi-polynomail in formula (13) resides only in the discrete moments
$\mu_{n}(\Pi_{v},t):=\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{n},$ (15)
for all $t>0$. We remark that prima facie these discrete moments may not appear to be polynomial-time computable. However, see Conjecture 1 below.
Remark 3.
We quickly recover a couple of the main results of Eva Linke’s seminal paper [28]. Linke’s important paper opened the door to the study of all positive real dilates of rational polytopes. Theorem 1.5 of [28] proves that the coefficients of quasi-polynomials of rational polytopes are always piecewise polynomial functions of $t$. Here it follows easily from Theorem 1. We may rewrite (15) as follows:
$\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k}=\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\langle q-tv,z\rangle}^{d-k}$ $=\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d-k},$
which is clearly a piecewise polynomial function of $t\in\mathbb{R}$, because the domain of summation is piecewise constant in $t$, while the summands are growing as a polynomial of degree dkd-k in $t$. Moreover, using the latter piecewise polynomial expression for the discrete moments of the parallelepiped, we may easily differentiate the $r$’th quasi-coefficient (14), in $t$:
$\frac{d}{dt}c_{r}(t)=\frac{d}{dt}\left(\frac{(-1)^{d}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d-r-j}\right)$
$=-\frac{(-1)^{d}}{d!}\sum_{v\in V}\langle v,z\rangle^{r+1}\sum_{j=0}^{d-r-1}\tfrac{d!(d-r-j)}{r!j!(d-r-j)!}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d-r-j-1}$
$=-(r+1)c_{r+1}(t),$
recovering one of the other main results in Linke’s paper, namely [28, Theorem 1.6].
Remark 4.
When the vertices are rational points $v\in\mathbb{Q}^{d}$, and we let the lattice flow proceed for all sufficiently large positive values of $t$, we topologically get a finite collection of closed geodesics on the torus, where each geodesic is an orbit of one integer point in $\mathcal{P}$}. As we see from Theorem 1, the periodicity of the Ehrhart quasi-polynomials arises precisely because of the periodicty of these closed geodesics.
An immediate corollary of Theorem 1 is the following expansion of the Ehrhart polynomial of a simple integer polytope $\mathcal{P}$}, in terms of the Barnes polynomials and of the discrete moments of its vertex parallelepipeds. Its proof is trivial: when all the vertices $v\in{\mathcal{P}}$ are integer points, and $t\in\mathbb{Z}$, then $\mathbb{Z}^{d}-tv=\mathbb{Z}^{d}$. It follows from (13) that there is no longer any periodic behavior in $L_{\mathcal{P}}(t)$.
Corollary 1.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a simple $d$-dimensional integer polytope. Then we have:
  1. (a)
    The Ehrhart polynomial of $\mathcal{P}$} is
    $L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{k=0}^{d}\binom{d}{k}\sum_{v\in V}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-k},$ (16)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t\in\mathbb{Z}_{>0}$.
  2. (b)
    The coefficient of $t^{r}$ in $L_{\mathcal{P}}(t):=\sum_{r=0}^{d}c_{r}t^{r}$ is equal to:
    $c_{r}=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-r-j},$ (17)
    valid for $t\in\mathbb{Z}_{>0}$, for almost all $z\in\mathbb{C}^{d}$, and for each $0\leq r\leq d$. The Barnes polynomials $B_{d}(x,{\mathbf{a}})$ are defined in (1), the Barnes numbers $B_{j}({\mathbf{a}})$ are defined in (8), and the vector of linear forms ${\mathbf{a}}_{v}$ is defined in (3). $\square$
Next, we obtain formulas for the discrete moments of positive dilates of a rational polytope $\mathcal{P}$}, in terms of the Barnes polynomials, and of the discrete moments of the fundamental parallelepipeds $\Pi_{v}$.
Theorem 2.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional, simple rational polytope.
  1. (a)
    For each positive integer $m$, we have the following formula for the $m$’th discrete moment of positive dilates of $\mathcal{P}$}:
    $\sum_{p\in t{\mathcal{P}}\cap\mathbb{Z}^{d}}\langle p,z\rangle^{m}$ $=(-1)^{d}m!\sum_{v\in V}\sum_{k=0}^{d+m}\tfrac{1}{k!(d+m-k)!}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-k},$ (18)
    a quasi-polynomial, valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
  2. (b)
    The coefficient of $t^{r}$ in the quasi-polynomial $\sum_{p\in t{\mathcal{P}}\cap\mathbb{Z}^{d}}\langle p,z\rangle^{m}:=\sum_{r=0}^{d}d_{r}(t)t^{r}$ of part (a) equals:
    $d_{r}(t)=\tfrac{(-1)^{d}m!}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-r-j},$ (19)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
  3. (c)
    The quasi-polynomial coefficients $d_{r}(t)$ above satisfy the following ordinary differential equation:
    $\frac{d}{dt}d_{r+1}(t)=-(r+1)d_{r+1}(t).$ (20)
(Proof) $\square$
Remark 5.
The ODE (20) in Theorem 2, part (c), directly extends Linke’s ODE in her seminal paper [28]. It is amusing - and a bit surprising - that we obtain exactly the same ODE, even though our ODE describes discrete moments of dilated rational polytopes.
The following collection of vanishing identities, arising from the proof of Theorem 1, give us information about the location of points in simple integer polytopes. Hence the following identities might be helpful in integer linear programming. From a computational perspective, these vanishing identities allow us to relate discrete moments of parallelepipeds recursively to each other, using lower discrete moments.
Theorem 3.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a simple $d$-dimensional integer polytope. Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$:
  1. (a)
    For each $0\leq m\leq d-1$, we have:
    $0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}.$ (21)
  2. (b)
    Equivalently, we may also express the latter set of identities by using the Barnes numbers, and removing the variable $t$. For each $(m,r)$ with $0\leq m\leq d-1,\text{ and }0\leq r\leq m$, we have:
    $0$ $=\sum_{j=0}^{m-r}\binom{m-r}{j}\sum_{v\in\rm{V}}\langle v,z\rangle^{r}B_{j}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-r-j}.$ (22)
    (Proof) $\square$
For example, a special case of Theorem 3, for $m=1,r=0$, as we show below in (106), is:
$\sum_{v\in\rm{V}}\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=\frac{1}{2}\sum_{v\in\rm{V}}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle},$ (23)
allowing us to rapidly compute certain weighted linear combinations of discrete moments of fundamental parallelepipeds, over all vertices of $\mathcal{P}$}. The cases with $r=m$ appeared in [12], while all the other cases of Theorem 3appear to be new. We expand on these special cases in (101) below. See also [37, Theorem 5.11] for a Fourier-based proof of these special cases.
A unimodular polytope is a simple, integer $d$-dimensional polytope ${\mathcal{P}}\subset\mathbb{R}^{d}$ such that at each vertex $v\in{\mathcal{P}}$, there is a collection of edge vectors $w_{1}(v),\dots w_{d}(v)$ that forms a basis for $\mathbb{Z}^{d}$. Unimodular polytopes are sometimes called “smooth polytopes” in toric geometry, and “delzant polytopes” in symplectic geometry. It is well known that in practice, there are numerous results - regarding the discrete geometry of unimodular polytopes - which are easy to prove. From the present perspective, this ease is due to the fact that most of the discrete moments at their vertices vanish, as we observe next, leading to an extreme simplification of the main results.
Corollary 2.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional unimodular polytope. Then we have:
$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{d}(t\langle v,z\rangle,{\mathbf{a}}_{v}),$
for all $t\in\mathbb{Z}_{>0}$, and almost all $z\in\mathbb{C}^{d}$. Equivalently, we have:
$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\left(\binom{d}{j}\sum_{v\in V}B_{j}({\mathbf{a}})\langle v,z\rangle^{d-j}\right)t^{d-j},$ (24)
(Proof) $\square$
Now we consider ${\mathcal{P}}^{o}$, the interior of the polytope $\mathcal{P}$}. Ehrhart’s reciprocity law, namely that $L_{{\mathcal{P}}^{o}}(t)=(-1)^{d}L_{{\mathcal{P}}}(-t)$, gives us the following immediate corollary of Theorem 1 for open, rational simple polytopes. All of the results above easily carry over to open polytopes, with their open vertex tangent cones. The only difference is the consideration of the interior of a fundamental parallelepiped, which is defined by:
$\text{int}(\Pi_{v}):=\{\sum_{k=1}^{d}\lambda_{k}w_{k}(v)\mid 0<\lambda<1\},$ (25)
for each vertex $v\in{\mathcal{P}}$. In all of the corollaries, we keep using the same notation as above for $B_{d}(x,{\mathbf{a}})$, $B_{j}({\mathbf{a}})$, and ${\mathbf{a}}_{v}$.
Corollary 3.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional simple rational polytope. Let ${\mathcal{P}}^{o}$ be the corresponding open polytope, and let $\text{int}(\Pi)$ be the open parallelepiped at the vertex $v\in{\mathcal{P}}$.
  1. (a)
    The quasi-polynomial of ${\mathcal{P}}^{o}$ is equal to:
    $L_{{\mathcal{P}}^{o}}(t)=\frac{1}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-t\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\text{int}(\Pi_{v}))}{\langle p,z\rangle}^{d-k},$ (26)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
  2. (b)
    The coefficient of $t^{r}$ in $L_{{\mathcal{P}}^{o}}(t):=\sum_{r=0}^{d}b_{r}(t)t^{r}$ is equal to:
    $b_{r}(t)=\frac{(-1)^{r}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\text{int}(\Pi_{v}))}{\langle p,z\rangle}^{d-r-j},$ (27)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$. $\square$
Remark 6.
We notice that the only difference between the periodicity of the quasi-coefficients for the interior of $\mathcal{P}$}, versus the corresponding periodicity of the quasi-coefficients for $\mathcal{P}$}, is that for each vertex the lattice flow on the torus (defined in (11) above) is now flowing in the opposite direction.
A lattice polytope is called hollow if its interior contains no integer points. There is considerable current research on classifying hollow simplices [1]. An immediate consequence of the latter corollary is the following equivalence for this property.
Corollary 4.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional integer, simple polytope. Then $\mathcal{P}$} is hollow if and only if
$0=\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\Pi_{v}\cap Z^{d}}{\langle p,z\rangle}^{d-k},$ (28)
for almost all $z\in\mathbb{C}^{d}$. $\square$
We also get the following interesting identities that arise from the standard topological fact that the constant term of the Ehrhart polynomial of an integer polytope is always $1$.
Corollary 5.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional integer, simple polytope. Then we have:
$1=\frac{(-1)^{d}}{d!}\sum_{k=0}^{d}\binom{d}{k}\sum_{v\in V}B_{k}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-k},$ (29)
valid for almost all $z\in\mathbb{C}^{d}$. $\square$
Proof.
To get the constant term of the Ehrhart polynomial, we simply substitute $t=0$ in Corollary 1 to obtain:
$1=L_{\mathcal{P}}(0)=\frac{(-1)^{d}}{d!}\sum_{k=0}^{d}\binom{d}{k}\sum_{v\in V}B_{k}(0,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-k},$ (30)
and the result follows because $B_{k}(0,{\mathbf{a}}_{v}):=B_{k}({\mathbf{a}}_{v})$, the Barnes numbers. ∎
Remark 7.
Specializng Corollary 5 to unimodular polytopes, we have:
$1=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{k}({\mathbf{a}}_{v}),$ (31)
and the proof is identical to the proof of Corollary 2. For example, it follows from (31) that for integer unimodular polygons, we have the amusing identity:
$12=\sum_{v\in\rm{V}}\left(\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+3\right),$ (32)
valid for almost all $z\in\mathbb{C}^{2}$ (see Example 5). Such identities appear to be new. The integer $12$ is easy to explain - it arises naturally from the second classical Bernoulli number $B_{2}=\frac{1}{6}$. Similarly, we have a corresponding identity for all $3$-dimensional unimodular polytopes:
$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{1}{2}\right).$
There is now a vast literature on Ehrhart polynomials and quasi-polynomials, and the reader may consult, for example, the books [4] [9] [21] [25] [37] [38], and papers [2] [5] [10] [12] [14] [15] [16] [17] [18] [19] [20] [22] [23] [24] [26] [29] [30] [31] [33] [34] [35] [39]. We remark that given a triangulation of any rational polytope into partially-open simplices, one can obtain similar formulas as in this paper. However, for general rational polytopes, finding an explicit triangulation is computationally expensive, even for general $3$-dimensional polytopes: some of its vertices might have $2^{100}$ edges, for example.
2 Known properties of Barnes polynomials
Given their defining generating function (1), it is natural that the Barnes polynomials should inherit some of the properties of the classical Bernoulli polynomials. First, the Barnes polynomials can be easily given in terms of a finite sum of classical Bernoulli polynomials:
$\frac{1}{k!}B_{k}(t,{\mathbf{a}})=\sum_{m_{1}+\cdots+m_{d}=k}\frac{a_{1}^{m_{1}-1}}{m_{1}!}\cdots\frac{a_{d}^{m_{d}-1}}{m_{d}!}B_{m_{1}}\Big(\tfrac{t}{a_{1}+\cdots+a_{d}}\Big)\cdots B_{m_{d}}\Big(\tfrac{t}{a_{1}+\cdots+a_{d}}\Big),$ (33)
where the $B_{m}(x)$ are the classical Bernoulli polynonomials, ${\mathbf{a}}:=(a_{1},\dots,a_{d})\in\mathbb{C}^{d}$, and the indices are nonnegative integers $m_{j}\geq 0$ (see Bayad and Kim [7, Theorem 2]). We see that if all of the $a_{j}$’s are rational numbers, then $B_{k}(t,{\mathbf{a}})$ is a polynomial in $t$, with rational coefficients. From (9) and (10), it follows at once that $(a_{1}a_{2}\cdots a_{d})B_{k}(t,{\mathbf{a}})\in\mathbb{Q}[a_{1},\dots,a_{d}]$. Also, it’s clear that the coefficients of $B_{k}(t,{\mathbf{a}})$, as a polynomial in $t$, are symmetric functions of $a_{1},\dots,a_{d}$. The Barnes polynomials enjoy the following properties.
[reciprocity law]
$(-1)^{k}B_{k}(-t,{\mathbf{a}})=B_{k}(t+a_{1}+\cdots+a_{d},{\mathbf{a}}),$ (34)
where ${\mathbf{a}}:=(a_{1},\dots,a_{d})^{T}$.
[homogeneity]
$B_{k}(\lambda t,\lambda{\mathbf{a}})=\lambda^{k-d}B_{k}(t,{\mathbf{a}}),$ (35)
for each $k\geq 1$, and for any fixed nonzero $\lambda\in\mathbb{C}$. Here by definition $\lambda{\mathbf{a}}:=(\lambda a_{1},\dots,\lambda a_{d})$. See [7] for proofs and more background on Barnes polynomials.
Using identity (9), we list the first few Barnes numbers with ${\mathbf{a}}\in\mathbb{C}^{d}$:
$B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}.$ (36)
$B_{1}({\mathbf{a}})=-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (37)
$(a_{1}\cdots a_{d})B_{2}({\mathbf{a}})=\frac{1}{6}\sum_{k=1}^{d}a_{k}^{2}+\frac{1}{2}\sum_{1\leq i<j\leq d}a_{i}a_{j}.$ (38)
$(a_{1}\cdots a_{d})B_{3}({\mathbf{a}})=-\frac{1}{4}\sum_{1\leq i\not=j\leq d}a_{i}^{2}a_{j}-\frac{1}{8}\sum_{1\leq i<j<k\leq d}a_{i}a_{j}a_{k}.$ (39)
In addition, using the identities (10) and (9), we list the first few Barnes polynomials with ${\mathbf{a}}\in\mathbb{C}^{d}$:
$B_{0}(t,{\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}},$ (40)
$B_{1}(t,{\mathbf{a}})$ $=\binom{1}{0}B_{0}({\mathbf{a}})t+\binom{1}{1}B_{1}({\mathbf{a}})$ (41)
$=\frac{1}{a_{1}\cdots a_{d}}t-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (42)
$B_{2}(t,{\mathbf{a}})=\binom{2}{0}B_{0}({\mathbf{a}})t^{2}+\binom{2}{1}B_{1}({\mathbf{a}})t+\binom{2}{2}B_{2}({\mathbf{a}})$ (43)
$=\frac{1}{a_{1}\cdots a_{d}}t^{2}-\Big(\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}\Big)t+\Big(\frac{1}{6}\sum_{k=1}^{d}\frac{a_{k}^{2}}{a_{1}\cdots a_{d}}+\frac{1}{2}\sum_{1\leq i<j\leq d}\frac{a_{i}a_{j}}{a_{1}\cdots a_{d}}\Big).$ (44)
Finally, for the purposes of explicitly computing some Ehrhart polynomials, we list the first few relevant Barnes polynomials, for each dimension $d=1,2,3$, derived using the relations (9), (10), (36), (37), (38), and (39):
$B_{1}(t,(a_{1}))=\frac{1}{a_{1}}t-\frac{1}{2}.$ (45)
$B_{2}\left(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right)\right)=t^{2}\left(\frac{1}{a_{1}a_{2}}\right)-t\left(\frac{1}{a_{2}}+\frac{1}{a_{1}}\right)+\frac{1}{6}\left(\frac{a_{2}}{a_{1}}+\frac{a_{1}}{a_{2}}\right)+\frac{1}{2}.$ (46)
$B_{3}\left(t,\left(\begin{smallmatrix}a_{1}\\ \\ a_{2}\\ \\ a_{3}\end{smallmatrix}\right)\right)=\binom{3}{0}B_{0}({\mathbf{a}})t^{3}+\binom{3}{1}B_{1}({\mathbf{a}})t^{2}+\binom{3}{2}B_{2}({\mathbf{a}})t+\binom{3}{3}B_{3}({\mathbf{a}})$ (47)
$=\frac{1}{a_{1}a_{2}a_{3}}t^{3}-\frac{3}{2}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}}\right)t^{2}$ (48)
$+\frac{3}{a_{1}a_{2}a_{3}}\left(\frac{1}{6}(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})+\frac{1}{2}(a_{1}a_{2}+a_{1}a_{3}+a_{2}a_{3})\right)t$ (49)
$+\frac{1}{a_{1}a_{2}a_{3}}\Big(-\frac{1}{4}\sum_{1\leq i\not=j\leq 3}a_{i}^{2}a_{j}-\frac{1}{8}\sum_{1\leq i<j<k\leq 3}a_{i}a_{j}a_{k}\Big)$ (50)
$=\frac{1}{a_{1}a_{2}a_{3}}t^{3}-\frac{3}{2}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}}\right)t^{2}$ (51)
$+\frac{1}{2}\left(\left(\frac{a_{1}}{a_{2}a_{3}}+\frac{a_{2}}{a_{1}a_{3}}+\frac{a_{3}}{a_{1}a_{2}}\right)+3\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\right)t$ (52)
$-\frac{1}{4}\left(\frac{a_{1}}{a_{2}}+\frac{a_{1}}{a_{3}}+\frac{a_{2}}{a_{1}}+\frac{a_{2}}{a_{3}}+\frac{a_{3}}{a_{1}}+\frac{a_{3}}{a_{2}}\right)-\frac{1}{8}.$ (53)
3 Dimension 1: intervals
It is instructive to consider the obligatory $1$-dimensional examples.
Example 1.
Consider the $1$-dimensional integer polytope ${\mathcal{P}}:=[a,b]$, where $a,b\in\mathbb{Z}$. Here the fundamental parallelepipeds at both vertex $a$ and vertex $b$ are unit intervals. The relevant Barnes polynomials in dimensions $1$ are: (see (40) and (45)):
$B_{0}(t,a_{1})=\frac{1}{a_{1}},\quad B_{1}(t,a_{1})=\frac{1}{a_{1}}t-\frac{1}{2},$ (54)
and here the vector of linear forms is just ${\mathbf{a}}_{v}=z$ for the vertex $v:=a$, and ${\mathbf{a}}_{v}=-z$ for the vertex $v:=b$. In dimension $1$ all polytopes are unimodular, so by Corollary 2 we have:
$L_{\mathcal{P}}(t)$ $=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)=-1\cdot B_{1}(taz,z)-1\cdot B_{1}(tbz,-z)$
$=-\Big(\frac{taz}{z}-\frac{1}{2}+\frac{tbz}{-z}-\frac{1}{2}\Big)=t(b-a)+1,$
recovering the simplest $1$-dimensional case. $\square$
Example 2.
More generally, consider the $1$-dimensional real polytope ${\mathcal{P}}:=[\alpha,\beta]$, where $\alpha,\beta\in\mathbb{R}$. Here the fundamental parallelepipeds are still unit intervals, namely $\Pi_{\alpha}=[0,1]$, and $\Pi_{\beta}=[-1,0]$. By Theorem 1 we have the quasi-polynomial:
$\left|t{\mathcal{P}}\cap\mathbb{Z}^{d}\right|$ $=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k}$
$=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)-\sum_{v\in\rm{V}}B_{0}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}$
$=-\Big(\frac{t\alpha z}{z}-\frac{1}{2}+\frac{t\beta z}{-z}-\frac{1}{2}\Big)-\Big(\frac{1}{z}\{-t\alpha\}z+\frac{1}{-z}(\{-t\beta\}-1)z\Big)$
$=t(\beta-\alpha)+\{-t\beta\}-\{-t\alpha\}+1$
$=t(\beta-\alpha)+(1-\{t\beta\})-(1-\{t\alpha\})+1$
$=t(\beta-\alpha)+\{t\alpha\}-\{t\beta\}+1,$
recovering this well-known, easy result. $\square$
4 Dimension 2: polygons
Suppose we are given any integer polygon $\mathcal{P}$}, defined as the convex hull of its vertices $\{v_{1},\cdots v_{N}\}$. At each vertex $v$, we are given its edge vectors $w_{1}(v),w_{2}(v)$. In the ensuing formula, we define $c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v))$, and the indicator function $1_{\mathbb{Z}^{2}}(x):=\begin{cases}1&\text{if $x\in\mathbb{Z}^{2}$}\\ 0&\text{otherwise}.\end{cases}$
An explicit description for the Ehrhart polynomial of $\mathcal{P}$}, using Corollary 1, is as follows:
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+\left(\frac{1}{2}\sum_{v\in V}\langle v,z\rangle\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)\right)t+1,$ (55)
valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$. To show how this follows from Theorem 1, we first recall the definition ${\mathbf{a}}_{v}:=\left(\begin{smallmatrix}\langle w_{1}(v),z\rangle\\ \langle w_{2}(v),z\rangle\end{smallmatrix}\right)$ at each vertex $v$ of $\mathcal{P}$}. From Theorem 1, we have:
$L_{\mathcal{P}}(t)$ $=\frac{1}{2}\sum_{k=0}^{2}\binom{2}{k}\sum_{v\in V}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}^{2-k}$
$=\frac{1}{2}\sum_{v\in V}B_{0}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}^{2}+\sum_{v\in V}B_{1}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}$
$+\frac{1}{2}\sum_{v\in V}B_{2}(t\langle v,z\rangle,{\mathbf{a}}_{v})\operatorname{vol}\Pi_{v}$
$=\frac{1}{2}\sum_{v\in V}\frac{\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}^{2}}{\langle{\mathbf{w}}_{1}(v),z\rangle,\langle{\mathbf{w}}_{2}(v),z\rangle}$
$+\sum_{v\in V}\left(\frac{1}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}t\langle v,z\rangle-\frac{1}{2}\frac{\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\right)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}$
$+\frac{1}{2}\sum_{v\in V}\left(t^{2}\langle v,z\rangle^{2}\left(\frac{\operatorname{vol}\Pi_{v}}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\right)-t\langle v,z\rangle\operatorname{vol}\Pi_{v}\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\right$.
$\left.\quad\quad+\frac{1}{6}\operatorname{vol}\Pi_{v}\left(\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}\right)+\frac{1}{2}\operatorname{vol}\Pi_{v}\right).$
To summarize, we have:
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+t\sum_{v\in V}\frac{\langle v,z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\left({\displaystyle\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}}-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle\right)+1,$ (56)
valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$. We used Corollary 5 for the constant term in (56), and equation (111) for the area of the polygon. Recalling that $c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v))$, the center of mass of the closure of the parallelogram $\Pi_{v}$, it’s an easy exercise that summing all the points in a fundamental parallelepiped gives us:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}p=\big(\operatorname{vol}\Pi_{v}+1-1_{\mathbb{Z}^{2}}(c_{v})\big)\ c_{v}.$ (57)
(hint: translate $\Pi_{v}\cap\mathbb{Z}^{2}$ to the origin by moving its center of mass, and notice that the new non-zero points cancel in paris). So we have:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}\langle p,z\rangle-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle=(1-1_{\mathbb{Z}^{2}}(c_{v}))\langle c_{v},z\rangle.$
Substituting the latter expression into (56), we obtain
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+t\sum_{v\in V}\frac{\langle v,z\rangle\langle c_{v},z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)+1$
$={\rm area}({\mathcal{P}})t^{2}+\left(\frac{1}{2}\sum_{v\in V}\langle v,z\rangle\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)\right)t+1,$
as claimed.
Example 3.
Consider the rational triangle $\mathcal{P}$} defined by ${\mathcal{P}}:=\operatorname*{conv}\left\{\left(\begin{smallmatrix}-\frac{1}{2}\\ -\frac{1}{4}\end{smallmatrix}\right),\left(\begin{smallmatrix}\frac{7}{2}\\ \frac{3}{4}\end{smallmatrix}\right),\left(\begin{smallmatrix}\frac{3}{2}\\ \frac{11}{4}\end{smallmatrix}\right)\right\}$, in Figure 1. In this example, we focus attention on its vertex $v=\left(\begin{smallmatrix}-\frac{1}{2}\\ -\frac{1}{4}\end{smallmatrix}\right)$, whose tangent cone $\mathcal{K}_{v}$ is also depicted in Figure 1.
The partial lattice flow shown on the left of Figure 2 is the union of the purple and red line segments, and is defined by
$\left\{LatticeFlow(t)\ \mid\ 0\leq t\leq\frac{1}{4}\right\}:=\left\{\mathbb{Z}^{2}-tv\pmod{\Pi_{v}}\ \big|\ 0\leq t\leq\frac{1}{4}.\right\}$
Each integer point is flowing along its geodesic defined on the flat $2$-torus given by $\Pi_{v}$. The lattice flow reaches its full closed geodesic flow, at $t=4$, as shown on the right-hand side of Figure 2. We note that there are exactly two closed geodesics here (one in red and the other in purple). Considering the integer points in $\Pi_{v}$ as a finite abelian group $G$ under addition mod $\Pi$, we see that the origin is contained in a subgroup of $G$ that has order $5$, and all closed geodesics arise from the cosets of this subgroup.
Figure 1: Top left: a rational triangle $\mathcal{P}$} with a rational vertex $v$. right: the vertex tangent cone $\mathcal{K}_{v}$ at the vertex $v$. Bottom left: the fundamental parallelepiped $\Pi_{v}$, with its integer edge vectors $w_{1}(v),w_{2}(v)$. Note that $\Pi_{v}$ always has a vertex at the origin.
Figure 2: Left: The fundamental parallelepiped $\Pi_{v}$ for the vertex $v:=\left(\begin{smallmatrix}-\frac{1}{2}\\ -\frac{1}{4}\end{smallmatrix}\right)$. The lattice flow $\{LatticeFlow(t)\ \mid\ 0\leq t\leq\frac{1}{4}\}$ is drawn on the left with red and purple line segments. Each integer point is flowing along its geodesic on the 22-torus, from $t=0$ until $t=\frac{1}{4}$. Right: the same lattice flow has reached its full closed geodesic flow, at $t=4$.
We may also observe that for any lattice flow, the orbit of any integer point in $\Pi_{v}$ produces a closed geodesic that has the same length as any other integer point. The reason for these equal length geodesics is easy, and arises from the finite abelian group $\mathbb{Z}^{d}/\Lambda$ structure, where $\Lambda$ is the integer sublattice generated by the vertices of $\Pi_{v}$. Thus, for example in Figure 2, the total length of the red geodesics equals the total length of the blue geodesics, because the blue geodesics arise from an index $2$ subgroup of $\mathbb{Z}^{2}/\Lambda$ here.
$\square$
Example 4.
In $\mathbb{R}^{2}$, just for fun let’s retrive the Ehrhart polynomial for the standard triangle ${\mathcal{P}}:=\operatorname*{conv}\{\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\}$, using Barnes polynomials. We’ll use Corollary 2 of Theorem 1, since the standard triangle is a unimodular polygon. Recalling that the only Barnes polynomials that occur in dimension $2$ are:
$B_{0}(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right))=\frac{1}{a_{1}a_{2}},\quad\quad B_{1}(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right))=\frac{1}{a_{1}a_{2}}t-\frac{1}{2}\frac{a_{1}+a_{2}}{a_{1}a_{2}},$ (58)
and
$B_{2}\Big(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right)\Big)$ $=t^{2}\left(\frac{1}{a_{1}a_{2}}\right)-t\left(\frac{1}{a_{2}}+\frac{1}{a_{1}}\right)+\frac{1}{6}\left(\frac{a_{2}}{a_{1}}+\frac{a_{1}}{a_{2}}\right)+\frac{1}{2},$ (59)
by identity (46). We have the following edge vectors for the vertex tangent cones at each of the vertices $v_{1},v_{2},v_{3}$:
$v_{1}:=\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)\implies\omega_{1}(v_{1})=\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\omega_{2}(v_{1})=\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right).$
$v_{2}:=\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\implies\omega_{1}(v_{2})=\left(\begin{smallmatrix}-1\\ \ 1\end{smallmatrix}\right),\omega_{2}(v_{2})=\left(\begin{smallmatrix}-1\\ \ 0\end{smallmatrix}\right).$
$v_{3}:=\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\implies\omega_{1}(v_{3})=\left(\begin{smallmatrix}\ 1\\ -1\end{smallmatrix}\right),\omega_{2}(v_{3})=\left(\begin{smallmatrix}\ 0\\ -1\end{smallmatrix}\right).$
Hence by definition, we have ${\mathbf{a}}_{v_{1}}:=\Big(\langle z,\omega_{1}(v_{1})\rangle,\langle z,\omega_{2}(v_{1})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}\\ z_{2}\end{smallmatrix}\right)$. Similarly, by definition we have ${\mathbf{a}}_{v_{2}}:=\Big(\langle z,\omega_{1}(v_{2})\rangle,\langle z,\omega_{2}(v_{2})\rangle\Big)^{T}=\left(\begin{smallmatrix}-z_{1}+z_{2}\\ -z_{1}\end{smallmatrix}\right)$, and
${\mathbf{a}}_{v_{3}}:=\Big(\langle z,\omega_{1}(v_{3})\rangle,\langle z,\omega_{2}(v_{3})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}-z_{2}\\ -z_{2}\end{smallmatrix}\right)$. By Corollary 2 , we have
$2!L_{P}(t)=(-1)^{2}\Big(B_{2}(t\cdot 0,{\mathbf{a}}_{v_{1}})+B_{2}(tz_{1},{\mathbf{a}}_{v_{2}})+B_{2}(tz_{2},{\mathbf{a}}_{v_{3}})\Big)$
$=\frac{1}{6}\left(\tfrac{z_{1}}{z_{2}}+\tfrac{z_{2}}{z_{1}}\right)+\tfrac{1}{2}$
$+t^{2}z_{1}^{2}\left(\tfrac{1}{(-z_{1}+z_{2})(-z_{1})}\right)-tz_{1}\left(\tfrac{1}{-z_{1}+z_{2}}+\tfrac{1}{-z_{1}}\right)+\tfrac{1}{6}\left(\tfrac{-z_{1}+z_{2}}{-z_{1}}+\tfrac{-z_{1}}{-z_{1}+z_{2}}\right)+\tfrac{1}{2}$
$+t^{2}z_{2}^{2}\left(\tfrac{1}{(-z_{2}+z_{1})(-z_{2})}\right)-tz_{2}\left(\tfrac{1}{-z_{2}+z_{1}}+\tfrac{1}{-z_{2}}\right)+\tfrac{1}{6}\left(\tfrac{-z_{2}+z_{1}}{-z_{2}}+\tfrac{-z_{2}}{-z_{2}+z_{1}}\right)+\tfrac{1}{2}$
$=t^{2}+3t+2,$
and we see that all of the $z$’s miraculously cancel out, as predicted by Theorem 1. $\square$
5 The unit cube in dimension d
We consider the $d$-dimensional unit cube $\square:=[-1,1]^{d}$, another unimodular polytope. Although it is trivial to compute the Ehrhart polynomial in this case, namely $L_{\square}(t)=(2t+1)^{d}$, the verification will shed some light on our use of the Barnes polynomials, and even gives new identities (equation (61) below) for the discrete hypercube $\{-1,1\}^{d}$.
To begin, we label each of the $2^{d}$ vertices $v\in\square$ by associating to each vertex a choice of signs as follows:
$v:=\left(\begin{smallmatrix}\alpha_{1}\\ .\\ .\\ .\\ \alpha_{d}\end{smallmatrix}\right),$ (60)
where we fix each $\alpha_{k}\in\{-1,1\}$. It’s clear that for any such fixed vertex $v$, the primitive integer edge vectors that are incident with $v$ are: $w_{1}(v)=-\alpha_{1}{\mathbf{e}}_{1},\dots,w_{d}(v)=-\alpha_{d}{\mathbf{e}}_{d}$, where we are using the standard basis vectors ${\mathbf{e}}_{1},\dots,{\mathbf{e}}_{d}$. It follows that the vector of linear forms associated with the vertex $v$ is:
${\mathbf{a}}_{v}$ $:=\Big(\langle w_{1}(v),z\rangle,\cdots,\langle w_{d}(v),z\rangle\Big)^{T}$
$=\left(\langle-\alpha_{1}{\mathbf{e}}_{1},z\rangle,\cdots,\langle-\alpha_{d}{\mathbf{e}}_{d},z\rangle\right)^{T}$
$=-\left(\alpha_{1}z_{1},\cdots,\alpha_{d}z_{d}\right)^{T}.$
By Corollary 2, we have
$L_{\square}(t)$ $=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\sum_{v\in\rm{V}}\left(\operatorname{vol}\Pi_{v}\right)\binom{d}{j}B_{d-j}({\mathbf{a}}_{v})\langle v,z\rangle^{j}\ t^{j}$
$=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\sum_{(\alpha_{1},\cdots,\alpha_{d})\in\{-1,1\}^{d}}B_{d-j}({\mathbf{a}}_{v})(\alpha_{1}z_{1}+\cdots+\alpha_{d}z_{d})^{j}\binom{d}{j}t^{j},$
where we are summing over the finite hypercube $\{-1,1\}^{d}$, consisting of $2^{d}$ elements. Comparing this result with the known Ehrhart polynomial $L_{\square}(t)=(2t+1)^{d}=\sum_{j=0}^{d}2^{j}t^{j}\binom{d}{j}$, we observe the following identities for the Barnes numbers, and for the discrete hypercube:
$\frac{(-1)^{d}}{d!}\sum_{(\alpha_{1},\cdots,\alpha_{d})\in\{-1,1\}^{d}}(\alpha_{1}z_{1}+\cdots+\alpha_{d}z_{d})^{j}B_{d-j}\left(\begin{smallmatrix}-\alpha_{1}z_{1}\\ .\\ .\\ .\\ -\alpha_{d}z_{d}\end{smallmatrix}\right)=2^{j},$ (61)
for each $0\leq j\leq d$, and for almost all $z\in\mathbb{C}^{d}$.
We may notice that the case $j=0$ of identity (61) was already covered by Corollary 5, in the special case of unimodular polytopes. The case $j=d$ of identity (61) seems interesting, and using $B_{0}({\mathbf{a}}_{v})=\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}$, tells us that
$2^{d}$ $=\frac{(-1)^{d}}{d!}\sum_{(\alpha_{1},\cdots,\alpha_{d})\in\{-1,1\}^{d}}(\alpha_{1}z_{1}+\cdots+\alpha_{d}z_{d})^{d}B_{0}({\mathbf{a}}_{v})$ (62)
$=\frac{(-1)^{d}}{d!}\sum_{(\alpha_{1},\cdots,\alpha_{d})\in\{-1,1\}^{d}}\frac{(\alpha_{1}z_{1}+\cdots+\alpha_{d}z_{d})^{d}}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}$ (63)
$=\frac{(-1)^{d}}{d!}\sum_{(\alpha_{1},\cdots,\alpha_{d})\in\{-1,1\}^{d}}\frac{(\alpha_{1}z_{1}+\cdots+\alpha_{d}z_{d})^{d}}{\prod_{k=1}^{d}(-\alpha_{k}z_{k})}$ (64)
$=\frac{1}{d!}\sum_{(\alpha_{1},\cdots,\alpha_{d})\in\{-1,1\}^{d}}\frac{(\alpha_{1}z_{1}+\cdots+\alpha_{d}z_{d})^{d}}{\prod_{k=1}^{d}\alpha_{k}z_{k}},$ (65)
for almost all $z\in\mathbb{C}^{d}$. As we’ve seen above, we can also reverse-engineer our approach. Namely, if we have an integer polytope $\mathcal{P}$} for which we already know $L_{\mathcal{P}}(t)$, then the polytope $\mathcal{P}$} determines a set of identities for the Barnes numbers, as we saw in (61).
6 Proof of Theorem 1
When a vertex $v\in{\mathcal{P}}$ is an integer point, the identity $\sigma_{\Pi_{v}+v}(z)=e^{\langle v,z\rangle}\sigma_{\Pi_{v}}(z)$ holds. When $v\notin\mathbb{Z}^{d}$, however, the latter identity is no longer necessarily true. Therefore more care has to be taken in case we consider the fundamental parallelepiped of a rational vertex, or in case we replace $v$ with $tv$, for $t\in\mathbb{Q}$.
We also recall the standard fact that the integer point enumerator for the parallelepiped $\Pi_{v}+v$ is defined by
$\sigma_{\Pi_{v}+v}(z):=\sum_{p\in(\Pi_{v}+v)\cap\mathbb{Z}^{d}}e^{\langle p,z\rangle},$ (66)
for all $z\in\mathbb{C}^{d}$. Given a $d$-dimensional simplicial cone $\mathcal{K}_{v}\subset\mathbb{R}^{d}$, with any real apex $v\in\mathbb{R}^{d}$, and with $d$ linearly independent integer edge vectors $\omega_{1},\dots,\omega_{d}\in\mathbb{Z}^{d}$, we have:
$\sum_{n\in\mathcal{K}_{v}\cap\mathbb{Z}^{d}}e^{\langle n,z\rangle}=\frac{\sigma_{\Pi_{v}+v}(z)}{\prod_{k=1}^{d}\left(1-e^{\langle\omega_{k},z\rangle}\right)}.$ (67)
[9, Theorem 3.4]. We use the notation $\mathcal{K}_{v}-v:=\mathcal{K}_{0}(v)$, the translation of the cone $\mathcal{K}_{v}$ to the origin. We observe that $t\mathcal{K}_{0}(v)=\mathcal{K}_{0}(v)$ for all $t>0$, a trivial but useful fact. Hence, when we dilate $\mathcal{K}_{v}$ by any $t>0$, all of its edge vectors remain the same, and only its apex $v$ gets dilated:
$t\mathcal{K}_{v}=t(\mathcal{K}_{v}-v)+tv=t\mathcal{K}_{0}(v)+tv=\mathcal{K}_{0}(v)+tv.$ (68)
We denote the latter cone by $\mathcal{K}_{tv}:=\mathcal{K}_{0}(v)+tv$.
Proof.
(of Theorem 1)We begin with the discrete Brion Theorem:
$\sigma_{t{\mathcal{P}}}(z)=\sigma_{t\mathcal{K}_{v_{1}}}(z)+\cdots+\sigma_{t\mathcal{K}_{v_{N}}}(z),$ (69)
an identity that is valid for almost all $z\in\mathbb{C}$, and valid for any rational $t>0$. We will also use the observation (68) regarding the meaning of $\mathcal{K}_{tv}$ for all positive $t>0$.
Step $1$. From (68), it follows that the elementary identity (67) becomes:
$\sigma_{t\mathcal{K}_{v}}(z):=\sum_{n\in t\mathcal{K}_{v}\cap\mathbb{Z}^{d}}e^{\langle n,z\rangle}=\frac{\sigma_{\Pi_{v}+tv}(z)}{\prod_{k=1}^{d}\left(1-e^{\langle\omega_{k},z\rangle}\right)},$ (70)
for all $t>0$. We now substitute $z:=xz_{0}$, with $x>0$ and $z_{0}$ a generic direction (thus avoiding any singularities in the denominators), and we begin to expand $\sigma_{t\mathcal{K}_{v}}(xz_{0})$ as a Taylor series about $x=0$:
$\sigma_{t\mathcal{K}_{v}}(xz_{0})$ $=\frac{\sigma_{\Pi_{v}+tv}(xz_{0})}{\prod_{k=1}^{d}\Big(1-e^{x\langle z_{0},w_{k}(v)\rangle}\Big)}$ (71)
$=\frac{e^{-tx\langle v,z_{0}\rangle}\sigma_{\Pi_{v}+tv}(xz_{0})}{x^{d}}\frac{x^{d}e^{tx\langle v,z_{0}\rangle}}{\prod_{k=1}^{d}\Big(1-e^{x\langle z_{0},w_{k}(v)\rangle}\Big)}$ (72)
$=\frac{e^{-tx\langle v,z_{0}\rangle}\sigma_{\Pi_{v}+tv}(xz_{0})}{x^{d}}(-1)^{d}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle),{\mathbf{a}}_{v})\frac{x^{k}}{k!}$ (73)
valid for almost all $z_{0}\in\mathbb{C}^{d}$. In the penultimate step we used the definition of the Barnes polynomials (1):
$\frac{x^{d}e^{tx}}{(e^{a_{1}x}-1)\cdots(e^{a_{d}x}-1)}:=\sum_{k\geq 0}B_{k}(t,{\mathbf{a}})\frac{x^{k}}{k!},$ (74)
where we replaced $a_{k}$ by $\langle w_{k}(v),z_{0}\rangle$. In other words, we used the definition of the Barnes polynomial, but with the vector ${\mathbf{a}}_{v}:=\Big(\langle w_{1}(v),z_{0}\rangle,\dots,\langle w_{d}(v),z_{0}\rangle\Big)^{T}$. To recap, using (69), we may sum the latter contributions (73), one from each of the vertex tangent cones, to obtain:
$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=\sigma_{t\mathcal{K}_{v_{1}}}(xz_{0})+\cdots+\sigma_{t\mathcal{K}_{v_{N}}}(xz_{0})$
$=\sum_{v\in\rm{V}}e^{-tx\langle v,z_{0}\rangle}\sigma_{\Pi_{v}+tv}(xz_{0})(-1)^{d}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!},$ (75)
valid for almost all $z_{0}\in\mathbb{C}^{d}$.
Step $2$. Next, we massage a bit the contribution of $e^{-tx\langle v,z_{0}\rangle}\sigma_{\Pi_{v}+tv}(xz_{0})$ as well:
$e^{-tx\langle v,z_{0}\rangle}\sigma_{\Pi_{v}+tv}(xz_{0})$ $=e^{x\langle-tv,z_{0}\rangle}\sum_{p\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}e^{x\langle p,z_{0}\rangle}$
$=\sum_{p\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}e^{x\langle p-tv,z_{0}\rangle}$
$=\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}e^{x\langle q,z_{0}\rangle},$ (76)
where we have made the change of variable q:=ptvq:=p-tv in the last step above. This change of variable allows us to think of the integer point transform in (76) as a lattice flow on a torus, as defined above in (11). Namely, as the positive real $t$ increases, the points $q$ range over $\mathbb{Z}^{d}-tv\mod\Pi_{v}$, inside the fixed torus defined by the parallelepiped $\Pi_{v}$.
Inserting (76) into (6), we now have:
$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}e^{x\langle q,z_{0}\rangle}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (77)
$=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{n\geq 0}{\langle q,z_{0}\rangle}^{n}\frac{x^{n}}{n!}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (78)
$=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{n,k\geq 0}{\langle q,z_{0}\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}.$ (79)
Step $3$. We now compare the constant terms on both sides of (79). On its left-hand side, the constant term is $\sigma_{t{\mathcal{P}}}(0)=L_{\mathcal{P}}(t)$, by definition. On the right-hand side of (79), we need to collect all terms whose indices satisfy the additional constraint k+nd=0k+n-d=0:
$L_{\mathcal{P}}(t)$ $=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{k=0}^{d}\binom{d}{k}{\langle q,z_{0}\rangle}^{d-k}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})$ (80)
$=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d-k},$ (81)
valid for almost all $z_{0}\in\mathbb{C}^{d}$. This concludes the proof of part (a).
To prove part (b), we expand the Barnes polynomial in the latter formula, in terms of the Barnes numbers. Namely, we use $B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j}$, in equation (81):
$L_{\mathcal{P}}(t)$ $=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{k=0}^{d}\sum_{j=0}^{k}\binom{d}{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})t^{k-j}\langle v,z_{0}\rangle^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d-k},$ (82)
Finally, we just have to sum over all relevant $k,j$, such that kj:=rk-j:=r, with $r$ fixed, so that we may isolate the coefficient of $t^{r}$ in the counting function $|t{\mathcal{P}}\cap\mathbb{Z}^{d}|$. We use the following elementary combinatorial identity, when setting kj=rk-j=r, which holds for any function $F(j,k)$ on $\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}$:
$\sum_{k=0}^{d}\sum_{j=0}^{k}F(j,k)=\sum_{j=0}^{d-r}F(j,r+j).$
In our context, (82) therefore implies that the coefficient of $t^{r}$ in $L_{\mathcal{P}}(t)$ is:
$\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{j=0}^{d-r}\frac{d!(r+j)!}{(r+j)!(d-r-j)!j!r!}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d-k}$ (83)
$=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d-k},$ (84)
valid for almost all $z_{0}\in\mathbb{C}^{d}$. ∎
7 Proof of Theorem 2
We recall the simple observation that the integer point transform of $\mathcal{P}$} is a generating functions for the discrete moments over $\mathcal{P}$}. Precisely, we have:
$\sigma_{t{\mathcal{P}}}(xz_{0}):=\sum_{n\in t{\mathcal{P}}\cap\mathbb{Z}^{d}}e^{\langle n,xz_{0}\rangle}=\sum_{n\in t{\mathcal{P}}\cap\mathbb{Z}^{d}}\sum_{k\geq 0}\frac{x^{k}}{k!}{\langle n,z_{0}\rangle}^{k}:=\sum_{k\geq 0}\frac{x^{k}}{k!}\mu_{k}(t{\mathcal{P}},z_{0}).$ (85)
Proof.
(of Theorem 2)We begin with equation (79) from the proof of Theorem 1:
$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{n,k\geq 0}{\langle q,z_{0}\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}.$ (86)
We fix any integer $m>0$, and we compare the coefficients of $x^{m}$ on both sides of (86). Using (85), we trivially see that the coefficient of $x^{m}$ on the left-hand side of (86) is $\frac{1}{m!}\mu_{m}(t{\mathcal{P}},z_{0})$.
On the right-hand side of (86), we need to collect all terms whose indices satisfy the constraint k+nd=mk+n-d=m:
$\frac{1}{m!}\mu_{m}(t{\mathcal{P}},z_{0})$ $=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{k=0}^{d+m}\frac{1}{k!(d+m-k)!}{\langle q,z_{0}\rangle}^{d+m-k}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v}),$ (87)
valid for almost all $z_{0}\in\mathbb{C}^{d}$. This completes the proof of part (a).
Step $3$. To prove part (b), we expand the Barnes polynomial in the latter formula, in terms of the Barnes numbers. Namely, we use $B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j}$, in equation (87): (81):
$\frac{1}{m!}\mu_{m}(t{\mathcal{P}},z_{0})=(-1)^{d}\sum_{v\in V}\sum_{k=0}^{d+m}\sum_{j=0}^{k}\tfrac{1}{k!(d+m-k)!}\binom{k}{j}B_{j}({\mathbf{a}}_{v})t^{k-j}\langle v,z_{0}\rangle^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-k},$ (88)
Finally, we just have to sum over all relevant $k,j$, such that kj:=rk-j:=r, with $r$ fixed, so that we may isolate the coefficient of $t^{r}$ on the right-hand side. We use the following elementary combinatorial identity, when setting kj=rk-j=r, which holds for any function $F(j,k)$ on $\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}$:
$\sum_{k=0}^{d+m}\sum_{j=0}^{k}F(j,k)=\sum_{j=0}^{d+m-r}F(j,r+j).$
In our context, (88) therefore implies that the coefficient of $t^{r}$ equals:
$(-1)^{d}\sum_{v\in V}\sum_{j=0}^{d+m-r}\frac{1}{(d+m-r-j)!j!r!}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j}$ (89)
$=\frac{(-1)^{d}}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j},$ (90)
valid for almost all $z_{0}\in\mathbb{C}^{d}$.
To prove part (c), we simply differentiate the latter expression in $t$:
$\frac{d}{dt}d_{r+1}(t)=\frac{d}{dt}\left(\tfrac{(-1)^{d}m!}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-r-j}\right)$
$=(-1)^{d}m!\sum_{v\in V}\langle v,z\rangle^{r+1}\sum_{j=0}^{d-r-1}\tfrac{(d+m-r-j)}{r!j!(d+m-r-j)!}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d+m-r-j-1}$
$=-(r+1)d_{r+1}(t),$
8 Proof of Theorem 3: vanishing identities for integer simple polytopes
We may push the proof of Theorem 1 a bit further, to derive certain canonical vanishing identities for simple integer polytopes $\mathcal{P}$}, arising from the singular part of (79), which is a Laurent series in $x$. In all the results that follow, we will keep using the same notation. In particular, we recall the vector of linear forms, defined for each vertex $v$ of a simple integer polytope $\mathcal{P}$}, by
${\mathbf{a}}_{v}:=\begin{pmatrix}\langle w_{1}(v),z\rangle\\ \vdots\\ \langle w_{d}(v),z\rangle\end{pmatrix}$.
This vector of linear forms is the local geometric information that we require at each vertex $v$ of $\mathcal{P}$}. One important new feature of the following canonical identities is that they encode the location of integer points in the fundamental parallelepipeds $\Pi_{v}$ of each of the vertices of $\mathcal{P}$}.
Proof.
(of Theorem 3)Using equation (79) from the proof of Theorem 1, we recall:
$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in\rm{V}}\sigma_{\Pi_{v}}(xz_{0})\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}.$
To prove part (a), we begin by noting that the left-hand-side of the latter identity is an entire function of $x$. We therefore know that for each fixed $n<0$, the coefficient of any $x^{n}$ on the right-hand-side must be identically zero. We next expand the exponential function $e^{x\langle z_{0},p\rangle}$ as a Taylor series as well, obtaining:
$(-1)^{d}\sigma_{t{\mathcal{P}}}(xz_{0})=$ $\sum_{v\in V}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}e^{x\langle z_{0},p\rangle}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (91)
$=\sum_{v\in V}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\sum_{n\geq 0}{\langle z_{0},p\rangle}^{n}\frac{x^{n}}{n!}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (92)
$=\sum_{v\in V}\sum_{n,k\geq 0}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle z_{0},p\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}$ (93)
The left-hand side of (91) is an entire function of $x$. Therefore we cannot have any singularities in $x$, so that comparing the coefficient of $x^{m-d}$ on both sides, for each fixed value of $m\in\{0,\dots,d-1\}$, we must get vanishing identities. Precisely, we get the following $d$ identities for each integer polytope in $\mathbb{R}^{d}$, by summing over all $k,n\geq 0$ in (93) with $k+n=m$ fixed:
$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z_{0}\rangle}^{m-k}.$ (94)
To prove part (b), we simplify (94) further by noticing that we have an identity for all positive integer values of $t$. Therefore each power of $t$ must have a coefficient that vanishes identically. We may use the explicit polynomial representation of the Barnes-Bernoulli polynomials (10), namely $B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j}$, together with (94), implies that
$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in\rm{V}}\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{k-j}t^{k-j}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}$ (95)
$=\sum_{v\in\rm{V}}\sum_{k=0}^{m}\sum_{j=0}^{k}\binom{m}{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{k-j}t^{k-j}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}$ (96)
Now we just have to fix kj:=rk-j:=r, and set the coefficient of $t^{r}$ equal to zero. We may use the following elementary combinatorial identity, when setting kj=rk-j=r:
$\sum_{k=0}^{m}\sum_{j=0}^{k}F(j,k)=\sum_{j=0}^{m-r}F(j,r+j).$
In our context, (96) therefore implies, after a little simplification:
$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\frac{{\langle p,z\rangle}^{m-r-j}}{(m-r-j)!},$ (97)
for each $(m,r)$ with $0\leq m\leq d-1,\text{ and }0\leq r\leq m$. ∎
9 Proof of Corollary 2: Ehrhart for unimodular polytopes
Proof.
(of Corollary 2)The unimodularity assumption means that for all vertices $v\in{\mathcal{P}}$, there is exactly one integer point in the fundamental parallelepiped $\Pi_{v}$, namely the origin. Therefore the discrete moments, for each $0\leq k\leq d-1$, degenerate to:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-k}=0,$ (98)
for all vertices $v\in{\mathcal{P}}$. Moreover, for $k=d$, we have
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-k}=\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1=1,$
so that we have:
$L_{\mathcal{P}}(t)$ $=\frac{(-1)^{d}}{d!}\sum_{k=0}^{d}\binom{d}{k}\sum_{v\in V}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{d-k}$
$=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{d}(t\langle v,z\rangle,{\mathbf{a}}_{v}).$
10 Examples, and some identities for special cases of the Theorems
Example 5.
In dimension $d=2$, Corollary 5 had an immediate consequence, namely (31), which tells us that for any unimodular integer polygon $P\subset\mathbb{R}^{2}$ we have:
$1=\frac{1}{2}\sum_{v\in\rm{V}}B_{2}({\mathbf{a}}_{v})=\frac{1}{2}\sum_{v\in\rm{V}}\left(\frac{1}{6}\left(\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}\right)+\frac{1}{2}\right),$ (99)
for almost all $z\in\mathbb{C}^{2}$. In other words, we have the identity
$12=\sum_{v\in\rm{V}}\left(\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+3\right),$ (100)
valid for almost all $z\in\mathbb{C}^{2}$. $\square$
Example 6.
In dimension $d=3$, the same consequence (31) of Corollary 5 tells us that for any unimodular integer polytope $P\subset\mathbb{R}^{3}$, we have:
$1=\frac{(-1)^{3}}{3!}\sum_{v\in\rm{V}}B_{3}({\mathbf{a}}_{v}),$
and unpacking $B_{3}({\mathbf{a}}_{v})$ by using (53), we arrive at the identity
$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{1}{2}\right),$
valid for almost all $z\in\mathbb{C}^{3}$. $\square$
10.1 Special cases of Theorem 3
We expand on a few of the vanishing identities from Theorem 3. We first note that the particular cases $r=m$ recovers known identities, by Brion and Vergne [12] and that the other identities appear to be new. Precisely, for $r=m$ with $0\leq m\leq d-1$, these $d$ known identities are:
$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{m}B_{0}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1=\sum_{v\in V}\frac{\langle v,z\rangle^{m}\operatorname{vol}\Pi_{v}}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle},$ (101)
valid for almost all $z\in\mathbb{C}^{d}$. Here we used $B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}$, and $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1=|\Pi_{v}\cap\mathbb{Z}^{d}|=\operatorname{vol}\Pi_{v}$. The latter identity is a standard fact, which may be arrived using Fourier transforms of polytopes [12] [37, Theorem 5.11].
Regarding the new identities, we first discuss the case $m=1,r=0$, as it presents a special case that is of independent interest: equation (106) below.
Case $(m,r)=(1,0)$. Here Theorem 3 translates into:
$0$ $=\sum_{v\in\rm{V}}\sum_{j=0}^{1}\frac{B_{j}({\mathbf{a}}_{v})}{j!}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\frac{{\langle p,z\rangle}^{1-j}}{(1-j)!}$ (102)
$=\sum_{v\in\rm{V}}B_{0}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle+\sum_{v\in\rm{V}}B_{1}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1$ (103)
$=\sum_{v\in\rm{V}}\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle+\sum_{v\in\rm{V}}B_{1}({\mathbf{a}}_{v})\operatorname{vol}\Pi_{v}$ (104)
$=\sum_{v\in\rm{V}}\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle-\frac{1}{2}\sum_{v\in\rm{V}}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle},$ (105)
so that we have
$\sum_{v\in\rm{V}}\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=\frac{1}{2}\sum_{v\in\rm{V}}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}.$ (106)
Case (m,r)=(m,m1)(m,r)=(m,m-1), with $2\leq m\leq d$. Here Theorem 3 gives us
$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\sum_{j=0}^{1}\frac{B_{j}({\mathbf{a}}_{v})}{j!}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\frac{{\langle p,z\rangle}^{1-j}}{(1-j)!}$ (107)
$=\sum_{v\in\rm{V}}\frac{\langle v,z\rangle^{m-1}}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle-\frac{1}{2}\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}.$ (108)
In other words, we have
$\sum_{v\in\rm{V}}\frac{\langle v,z\rangle^{m-1}}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=\frac{1}{2}\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle},$ (109)
for all simple integer polytopes, for almost all $z\in\mathbb{C}^{d}$, and for each $1\leq m\leq d$.
Next, we make a simple observation that holds for the family of unimodular polytopes. If $\mathcal{P}$} is a unimodular polytope, then the latter identity (109) enjoys a simplification. For a unimodular polytope, we have $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=0$, because the origin is the only integer point in $\Pi_{v}$. Therefore
$\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}=0,$ (110)
for each fixed $1\leq m\leq d$, and for almost all $z\in\mathbb{C}^{d}$.
Remark 8.
Case $(m,r)=(d,d)$. An elementary fact from Ehrhart theory [37] is that the leading coefficient of $L_{\mathcal{P}}(t)$ is always equal to $\operatorname{vol}{\mathcal{P}}$. Recalling that $B_{0}({\mathbf{a}}_{v})=\frac{1}{a_{1}\cdots a_{d}}$, we see that
$\operatorname{vol}{\mathcal{P}}=\frac{(-1)^{d}}{d!}\sum_{v\in\rm{V}}\left(\operatorname{vol}\Pi_{v}\right)\langle v,z\rangle^{d}B_{0}({\mathbf{a}}_{v})=\frac{(-1)^{d}}{d!}\sum_{v\in\rm{V}}\left(\operatorname{vol}\Pi_{v}\right)\frac{\langle v,z\rangle^{d}}{\prod_{m=1}^{d}\langle w_{m}(v),z\rangle},$ (111)
for almost all $z\in\mathbb{C}^{2}$. We’ve recovered here a known volume identity of Brion [11, Corollary 2] and Lawrence [27] (see also [37, Theorem 7.23] for another proof).
11 Further remarks and research directions
We note that for fixed dimension $d$, the formulation for $L_{P}(t)$ in Corollary 1 is not a polynomial-time algorithm, because the finite sums defined by
$\mu_{n}(\Pi_{v}):=\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{n},$
at each vertex $v\in\Pi$, have exponentially many terms, as a function of the bit-input of $\Pi_{v}$. They are therefore not polynomial-time computable as they are defined. For simplicity, in this section we write the discrete moments for $t=1$ as $\mu_{n}(\Pi_{v})$.
Conjecture 1.
Fix the dimension $d$, and a half-open integer paralleleppiped $\Pi\subset\mathbb{R}^{d}$. Then for each fixed integer $k$, and fixed $z\in\mathbb{C}^{d}$, there is a polynomial-time algorithm, in the bit-input of the vertices of an integer parallelepiped $\Pi$, for computing the discrete moment $\mu_{k}(\Pi)$.
Remark 9.
We may give a trivial bound for the discrete moments $\mu_{n}(\Pi_{v})$ in Theorem 1, and hence for the coefficients of Ehrhart quasi-polynomials by considering $M_{v}:=\max_{p\in\Pi_{v}}\|p\|$. We have
$\|\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{n}\ \ \|\leq\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\|p\|^{n}\|z\|^{n}\leq M_{v}^{n}\|z\|^{n}\operatorname{vol}\Pi_{v}.$
Of course, the latter bound is far from optimal.
Question 1.
What is a good bound for the discrete moments $\mu_{n}(\Pi_{v})$?
There is a large literature regarding bounds of Ehrhart polynomials, see for example [18].
Remark 10.
Here is an idea to help us compute Ehrhart polynomials in practice, using Corollary 1. Suppose we have a particularly difficult vertex $v\in{\mathcal{P}}$, in the sense that its fundamental parallelepiped $\Pi_{v}$ has for example $2^{30}$ terms we need to sum. There is a way to omit this vertex and the corresponding sums $\mu_{k}(\Pi_{v})$ (for $k=1,2,\dots,d-1$) from our computations, but still compute $L_{\mathcal{P}}(t)$ precisely.
Let $d\geq 2$. We know from Section 2 that $B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\prod_{j=1}^{d}\langle w_{j},z\rangle$ is a nonconstant polynomial in $z_{1},\dots,z_{d}$, for all $k\geq 1$. Picking a generic $z\in\mathbb{C}^{d}$, in the sense of (7), already implies that $z\in\mathbb{C}^{d}$ satisfies $\prod_{j=1}^{d}\langle w_{j},z\rangle\not=0$. Now pick a generic $z\in\mathbb{C}^{d}$ that also belong to the variety defined by
$\bigcap_{k=1}^{d-1}\left\{z\in\mathbb{C}^{d}\ \mid\ B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v}).\prod_{j=1}^{d}\langle w_{j},z\rangle=0\right\}.$ (112)
This variety gives us extra flexibility to choose points $z\in\mathbb{C}^{d}$ that allow us to omit the discrete moment sums $\mu_{k}(\Pi_{v})$ (for $k=1,2,\dots,d-1$) from our calculations. Due to the flexibility in picking any generic $z\in\mathbb{C}^{d}$ in the theorems, we arrive at the following conclusion:
In computing the Ehrhart coefficients $c_{1},\dots c_{d-1}$ of Corollary 1, we may ignore any one fixed vertex of our choice.
Question 2.
Can we ignore two or more vertices, in computing the Ehrhart coefficients $c_{1},\dots c_{d-1}$ of Corollary 1, by using the variety in Remark 10?

### v2

@@
 From this work, it is clear that the complexity of computing Ehrhart quasi-polynomials lies mainly in the computation of various discrete moments of parallelepipeds.
-These discrete moments are in general summed over a particular lattice flow on a closed torus, defined in this paper.
-Some of the consequences involve novel vanishing identities for rational polytopes, novel formulations of Ehrhart polynomials of unimodular polytopes, and a differential equation that extends the work of Eva Linke.
+These discrete moments are in general summed over a particular lattice flow on a compact torus, defined in this paper.
+Some of the consequences involve novel vanishing identities for rational polytopes.
+As another consequence, we obtain a differential equation for discrete moments of rational polytopes, which extends the work of Eva Linke.
+For smooth polytopes, we obtain novel and much simpler formulations of Ehrhart polynomials, discrete moments, and vanishing identities that may be of independent interest from the perspective of Barnes polynomials and Barnes numbers.
+These formulations show the utility of Barnes polynomials in geometric combinatorics, due to their very rich structure that extends the $1$-dimensional Bernoulli polynomials.
 1 Introduction, and statements of the main results
@@
 [7, 8, 9].
-We use elementary techniques, and our first main result is an explicit formula for the Ehrhart polynomial (resp.
-quasi-polynomial) of any simple integer (resp.
-rational) polytope $\mathcal{P}$}, in terms of Barnes polynomials and discrete moments of parallelepipeds.
-In this sense, this paper also initiates the study of discrete moments of parallelepipeds and more generally the discrete moments of rational polytopes.
+We use elementary techniques, and our first main result is an explicit formula for the Ehrhart quasi-polynomial of any simple integer rational polytope $\mathcal{P}$}, in terms of Barnes polynomials and discrete moments of parallelepipeds.
+In this sense, this paper also initiates the study of discrete moments of half-open parallelepipeds and more generally the discrete moments of rational polytopes.
 Such discrete moments may be thought of as belonging to spatial statistics over the integer lattice [14].
-For our applications here, we note that if all of the $a_{k}$’s are rational numbers, then definition (1) gives us a polynomial $B_{k}(t,{\mathbf{a}})$ function of $t$, with rational coefficients (see (33) below).
 Throughout, we write $V$ for the set of vertices of a polytope $\mathcal{P}$}, and we define the usual local information at each vertex $v$ of a simple rational polytope $\mathcal{P}$}.
@@
 $\left\{z\in\mathbb{C}^{d}\mid\langle z,w(v)\rangle\notin 2\pi i\mathbb{Z},\text{ for any edge }w(v)\text{ of }{\mathcal{P}}\right\}.$ (7)
-Throughout, we will state our theorems with the words “for almost all $z\in\mathbb{C}^{d}$” to mean that they hold for all $z$ in the set (7) above.
+Throughout, we state our theorems with the words “for almost all $z\in\mathbb{C}^{d}$” to mean that they hold for all $z$ in the set (7) above.
 We also recall that a quasi-polynomial is a function $f(t)=\sum_{k=0}^{d}c_{k}(t)t^{k}$, where each coefficient $c_{k}(t)$ is a periodic function of $t\in\mathbb{Z}_{>0}$.
@@
 $B_{k}({\mathbf{a}})=\sum_{m_{1}+\cdots+m_{d}=k}\binom{k}{m_{1},\dots,m_{d}}a_{1}^{m_{1}-1}\cdots a_{d}^{m_{d}-1}B_{m_{1}}\cdots B_{m_{d}},$ (9)
-where the $B_{j}$’s are the classical Bernoulli numbers [7, Theorem 2]. In section 2 below, we give some additional background and examples of Barnes polynomials and Barnes numbers.
+where the $B_{j}$’s are the classical Bernoulli numbers [8, Theorem 2].
+Equation (9) also shows that $(a_{1}\cdots a_{d})B_{k}({\mathbf{a}})$ is a homogeneous polynomial of the $a_{j}$’s, of degree $k$.
+Appendix A gives some additional background and examples of Barnes polynomials and Barnes numbers.
 The Barnes numbers extend the usual Bernoulli numbers, in the precise sense that for dimension $d=1$, and for ${\mathbf{a}}=(1)$, we have $B_{j}({\mathbf{a}})=B_{j}$, the $j$’th Bernoulli number.
-A very useful explicit polynomial representation for the Barnes polynomials, in terms of the Barnes numbers $B_{k}({\mathbf{a}})$ is:
+In many of our results we will use the following explicit polynomial representation for the Barnes polynomials in terms of the Barnes numbers $B_{k}({\mathbf{a}})$:
 $B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j},$ (10)
@@
 $LatticeFlow(t):=\mathbb{Z}^{d}-tv\pmod{\Pi}$, (11)
-as in Figure 2, Section 4 below. We will compute in a straightforward way some explicit lattice flows over half-open parallelepipeds (see Sections 3 and 4).
+as in §2.2, Figure 2 below. We will compute in a straightforward way some explicit lattice flows over half-open parallelepipeds (see §2.1 and §2.2).
 Definition 1.
@@
 $\mu_{k}(\Pi,t):=\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi)}\langle p,z\rangle^{k},$ (12)
-for each fixed integer $k\geq 0$, $v\in\mathbb{R}^{d}$ and $t\geq 0$.
+for each fixed integer $k\geq 0$, $v\in\mathbb{R}^{d}$ and $t\in\mathbb{R}$.
 Theorem 1.
@@
 It may seem rather surprising a priori that the right-hand-sides of both (13) and (14) appears to depend on $z\in\mathbb{C}^{d}$; but in fact a posteriori they are independent of $z$, as the left-hand-side of (13) shows.
-Moreover, it is possible that judicious choices of $z$ may reduce the computational complexity of the ensuing discrete moments, and therefore of the Ehrhart polynomials and quasi-polynomials.
+Moreover, it is possible that judicious choices of $z$ may reduce the computational complexity of the ensuing discrete moments, and therefore of the Ehrhart polynomials and quasi-polynomials (see §8, Remark 10).
 Remark 2.
@@
 $\mu_{n}(\Pi_{v},t):=\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{n},$ (15)
-for all $t>0$. We remark that prima facie these discrete moments may not appear to be polynomial-time computable. However, see Conjecture 1 below.
+for all $t>0$. We remark that prima facie these discrete moments may not appear to be polynomial-time computable. However, Barvinok’s algorithm [6] shows that they are (see also Question 1 below).
 Remark 3.
-We quickly recover a couple of the main results of Eva Linke’s seminal paper [28].
-Linke’s important paper opened the door to the study of all positive real dilates of rational polytopes.
+We quickly recover a couple of the main results of Eva Linke’s important paper [32].
+Linke’s work opened the door to a more precise study of all positive real dilates of rational polytopes.
 Theorem 1.5 of [32] proves that the coefficients of quasi-polynomials of rational polytopes are always piecewise polynomial functions of $t$.
-Here it follows easily from Theorem 1.
-We may rewrite (15) as follows:
+Here it follows easily from Theorem 1 by rewriting (15) as follows:
 $\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k}=\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\langle q-tv,z\rangle}^{d-k}$ $=\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d-k},$
@@
 $=-(r+1)c_{r+1}(t),$
-recovering one of the other main results in Linke’s paper, namely [28, Theorem 1.6].
+recovering one of the other main results in Linke’s paper, namely [32, Theorem 1.6].
 Remark 4.
-When the vertices are rational points $v\in\mathbb{Q}^{d}$, and we let the lattice flow proceed for all sufficiently large positive values of $t$, we topologically get a finite collection of closed geodesics on the torus, where each geodesic is an orbit of one integer point in $\mathcal{P}$}.
-As we see from Theorem 1, the periodicity of the Ehrhart quasi-polynomials arises precisely because of the periodicty of these closed geodesics.
+Given a rational vertex $v\in\mathbb{Q}^{d}$, let’s allow the lattice flow $\mathbb{Z}^{d}-tv\pmod{\Pi}$ to proceed for all sufficiently large positive values of $t$.
+Topologically, we get a finite collection of closed geodesics on the torus, where each geodesic is an orbit of one integer point in $\mathcal{P}$}.
+As we see from Theorem 1, the periodicity of the Ehrhart quasi-polynomials arises precisely because of the periodicity of these closed geodesics, appearing in the discrete moments of the fundamental parallelepipeds of the vertices.
 An immediate corollary of Theorem 1 is the following expansion of the Ehrhart polynomial of a simple integer polytope $\mathcal{P}$}, in terms of the Barnes polynomials and of the discrete moments of its vertex parallelepipeds.
@@
 Next, we obtain formulas for the discrete moments of positive dilates of a rational polytope $\mathcal{P}$}, in terms of the Barnes polynomials, and of the discrete moments of the fundamental parallelepipeds $\Pi_{v}$.
+Remark 5.
+We observe that Theorem 1 also implies that for an integer polytope $\mathcal{P}$}, allowing all positive dilates $t>0$ quickly implies here that the periodic components always have a period of $1$, because for integral vertices $v\in{\mathcal{P}}$, the domain of summation in the discrete moments of formula (13) satisfies $\mathbb{Z}^{d}-v=\mathbb{Z}^{d}$.
+It becomes an interesting question of whether there exists a smaller period, $0<t<1$, for some integer polytopes.
 Theorem 2.
@@
     The quasi-polynomial coefficients $d_{r}(t)$ above satisfy the following ordinary differential equation:
-    $\frac{d}{dt}d_{r+1}(t)=-(r+1)d_{r+1}(t).$ (20)
+    $\frac{d}{dt}d_{r}(t)=-(r+1)d_{r+1}(t).$ (20)
 (Proof) $\square$
@@
 The ODE (20) in Theorem 2, part (c), directly extends Linke’s ODE in her seminal paper [32].
-It is amusing - and a bit surprising - that we obtain exactly the same ODE, even though our ODE describes discrete moments of dilated rational polytopes.
+It is amusing - and a bit surprising - that we obtain exactly the same ODE as in Linke’s work, even though our ODE describes discrete moments of dilated rational polytopes.
+1.1 Vanishing identities for rational polytopes
 The following collection of vanishing identities, arising from the proof of Theorem 1, give us information about the location of points in simple integer polytopes.
@@
 Theorem 3.
-Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a simple $d$-dimensional integer polytope.
-Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$:
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a simple $d$-dimensional rational polytope.
+Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$ and for all $t>0$:
   1. (a)
     For each $0\leq m\leq d-1$, we have:
-    $0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}.$ (21)
+    $0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{m-k}.$ (21)
   2. (b)
@@
 For each $(m,r)$ with $0\leq m\leq d-1,\text{ and }0\leq r\leq m$, we have:
-    $0$ $=\sum_{j=0}^{m-r}\binom{m-r}{j}\sum_{v\in\rm{V}}\langle v,z\rangle^{r}B_{j}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-r-j}.$ (22)
+    $0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!(m-r-j)!}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-r-j}.$ (22)
     (Proof) $\square$
-For example, a special case of Theorem 3, for $m=1,r=0$, as we show below in (106), is:
+For example, in the special case of integer polytopes, Theorem 3 with $m=1,r=0$ tells us that
 $\sum_{v\in\rm{V}}\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=\frac{1}{2}\sum_{v\in\rm{V}}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle},$ (23)
-allowing us to rapidly compute certain weighted linear combinations of discrete moments of fundamental parallelepipeds, over all vertices of $\mathcal{P}$}.
+as we show below in (94).
+This allows us to rapidly compute certain weighted linear combinations of discrete moments of fundamental parallelepipeds, over all vertices of $\mathcal{P}$}.
 The cases with $r=m$ appeared in [13], while all the other cases of Theorem 3 appear to be new.
-We expand on these special cases in (101) below.
-See also [37, Theorem 5.11] for a Fourier-based proof of these special cases.
-A unimodular polytope is a simple, integer $d$-dimensional polytope ${\mathcal{P}}\subset\mathbb{R}^{d}$ such that at each vertex $v\in{\mathcal{P}}$, there is a collection of edge vectors $w_{1}(v),\dots w_{d}(v)$ that forms a basis for $\mathbb{Z}^{d}$.
-Unimodular polytopes are sometimes called “smooth polytopes” in toric geometry, and “delzant polytopes” in symplectic geometry.
-It is well known that in practice, there are numerous results - regarding the discrete geometry of unimodular polytopes - which are easy to prove.
-From the present perspective, this ease is due to the fact that most of the discrete moments at their vertices vanish, as we observe next, leading to an extreme simplification of the main results.
-Corollary 2.
-Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional unimodular polytope. Then we have:
-$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{d}(t\langle v,z\rangle,{\mathbf{a}}_{v}),$
-for all $t\in\mathbb{Z}_{>0}$, and almost all $z\in\mathbb{C}^{d}$. Equivalently, we have:
-$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\left(\binom{d}{j}\sum_{v\in V}B_{j}({\mathbf{a}})\langle v,z\rangle^{d-j}\right)t^{d-j},$ (24)
-(Proof) $\square$
-Now we consider ${\mathcal{P}}^{o}$, the interior of the polytope $\mathcal{P}$}.
-Ehrhart’s reciprocity law, namely that $L_{{\mathcal{P}}^{o}}(t)=(-1)^{d}L_{{\mathcal{P}}}(-t)$, gives us the following immediate corollary of Theorem 1 for open, rational simple polytopes.
-All of the results above easily carry over to open polytopes, with their open vertex tangent cones.
-The only difference is the consideration of the interior of a fundamental parallelepiped, which is defined by:
-$\text{int}(\Pi_{v}):=\{\sum_{k=1}^{d}\lambda_{k}w_{k}(v)\mid 0<\lambda<1\},$ (25)
-for each vertex $v\in{\mathcal{P}}$. In all of the corollaries, we keep using the same notation as above for $B_{d}(x,{\mathbf{a}})$, $B_{j}({\mathbf{a}})$, and ${\mathbf{a}}_{v}$.
-Corollary 3.
-Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional simple rational polytope.
-Let ${\mathcal{P}}^{o}$ be the corresponding open polytope, and let $\text{int}(\Pi)$ be the open parallelepiped at the vertex $v\in{\mathcal{P}}$.
-  1. (a)
-    The quasi-polynomial of ${\mathcal{P}}^{o}$ is equal to:
-    $L_{{\mathcal{P}}^{o}}(t)=\frac{1}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-t\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\text{int}(\Pi_{v}))}{\langle p,z\rangle}^{d-k},$ (26)
-    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
-  2. (b)
-    The coefficient of $t^{r}$ in $L_{{\mathcal{P}}^{o}}(t):=\sum_{r=0}^{d}b_{r}(t)t^{r}$ is equal to:
-    $b_{r}(t)=\frac{(-1)^{r}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\text{int}(\Pi_{v}))}{\langle p,z\rangle}^{d-r-j},$ (27)
-    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$. $\square$
-Remark 6.
-We notice that the only difference between the periodicity of the quasi-coefficients for the interior of $\mathcal{P}$}, versus the corresponding periodicity of the quasi-coefficients for $\mathcal{P}$}, is that for each vertex the lattice flow on the torus (defined in (11) above) is now flowing in the opposite direction.
-A lattice polytope is called hollow if its interior contains no integer points.
-There is considerable current research on classifying hollow simplices [1].
-An immediate consequence of the latter corollary is the following equivalence for this property.
-Corollary 4.
-Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional integer, simple polytope. Then $\mathcal{P}$} is hollow if and only if
-$0=\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\Pi_{v}\cap Z^{d}}{\langle p,z\rangle}^{d-k},$ (28)
-for almost all $z\in\mathbb{C}^{d}$. $\square$
-We also get the following interesting identities that arise from the standard topological fact that the constant term of the Ehrhart polynomial of an integer polytope is always $1$.
+We expand on these special cases $r=m$ in (89) below (See also [42, Theorem 5.11] for a Fourier-based proof).
+1.2 The constant term 1 identity
+The following interesting identities arise from the standard topological fact that the constant term of the Ehrhart polynomial of an integer polytope is always $1$.
 Corollary 2.
@@
 and the result follows because $B_{k}(0,{\mathbf{a}}_{v}):=B_{k}({\mathbf{a}}_{v})$, the Barnes numbers. ∎
+1.3 Open polytopes
+Now we consider ${\mathcal{P}}^{o}$, the interior of the polytope $\mathcal{P}$}.
+Ehrhart’s reciprocity law, namely that $L_{{\mathcal{P}}^{o}}(t)=(-1)^{d}L_{{\mathcal{P}}}(-t)$, gives us the following immediate corollary of Theorem 1 for open, rational simple polytopes.
+All of the results above easily carry over to open polytopes, with their open vertex tangent cones.
+In all of the corollaries, we keep using the same notation as above for $B_{d}(x,{\mathbf{a}})$, $B_{j}({\mathbf{a}})$, and ${\mathbf{a}}_{v}$.
+Corollary 3.
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional simple rational polytope. Let ${\mathcal{P}}^{o}$ be the corresponding open polytope.
+  1. (a)
+    The quasi-polynomial of ${\mathcal{P}}^{o}$ is equal to:
+    $L_{{\mathcal{P}}^{o}}(t)=\frac{1}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-t\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k},$ (26)
+    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
+  2. (b)
+    The coefficient of $t^{r}$ in $L_{{\mathcal{P}}^{o}}(t):=\sum_{r=0}^{d}b_{r}(t)t^{r}$ is equal to:
+    $b_{r}(t)=\frac{(-1)^{r}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-r-j},$ (27)
+    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$. $\square$
 Remark 7.
-Specializng Corollary 5 to unimodular polytopes, we have:
-$1=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{k}({\mathbf{a}}_{v}),$ (31)
-and the proof is identical to the proof of Corollary 2. For example, it follows from (31) that for integer unimodular polygons, we have the amusing identity:
+We notice that the only difference between the periodicity of the quasi-coefficients for the interior of $\mathcal{P}$}, versus the corresponding periodicity of the quasi-coefficients for $\mathcal{P}$}, is that for each vertex the lattice flow on the torus (defined in (11) above) is now flowing in the opposite direction.
+A lattice polytope is called hollow if its interior contains no integer points.
+There is considerable current research on classifying hollow simplices [1].
+An immediate consequence of the Corollary 3 is the following equivalence for this property.
+Corollary 4.
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional integer, simple polytope. Then $\mathcal{P}$} is hollow if and only if
+$0=\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\Pi_{v}\cap Z^{d}}{\langle p,z\rangle}^{d-k},$ (28)
+for almost all $z\in\mathbb{C}^{d}$. $\square$
+1.4 Smooth polytopes
+A smooth polytope is a simple, integer $d$-dimensional polytope ${\mathcal{P}}\subset\mathbb{R}^{d}$ such that at each vertex $v\in{\mathcal{P}}$, there is a collection of edge vectors $w_{1}(v),\dots w_{d}(v)$ that forms a basis for $\mathbb{Z}^{d}$.
+The terminology of Smooth polytopes comes from toric geometry, and smooth polytopes are sometimes also called “totally unimodular polytopes” in integer programming, or “delzant polytopes” in symplectic geometry.
+It is well known that in practice, there are numerous results - regarding the discrete geometry of smooth polytopes - which are easy to prove.
+From the present perspective, this ease is due to the fact that most of the discrete moments at their vertices vanish, as we observe next, leading to an extreme simplification of the main results.
+First, specializing Theorem 1 to smooth polytopes immediately gives us the following formula for their Ehrhart polynomials.
+The proofs of the rest of the corollaries are omitted, because their immediate proofs are identical to that of Corollary 5.
+Corollary 5.
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope. Then we have:
+$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{d}(t\langle v,z\rangle,{\mathbf{a}}_{v}),$
+for all $t\in\mathbb{Z}_{>0}$, and almost all $z\in\mathbb{C}^{d}$. Equivalently, we have:
+$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\left(\binom{d}{j}\sum_{v\in V}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{d-j}\right)t^{d-j},$ (29)
+for all $t\in\mathbb{Z}_{>0}$, and almost all $z\in\mathbb{C}^{d}$. (Proof) $\square$
+Specializing Theorem 2 to smooth polytopes gives the following moment formulas for smooth polytopes.
+Corollary 6.
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope.
+For each positive integer $m$, we have the following formula for the $m$’th discrete moment of positive dilates of $\mathcal{P}$}:
+$\sum_{p\in t{\mathcal{P}}\cap\mathbb{Z}^{d}}\langle p,z\rangle^{m}$ $=(-1)^{d}\frac{m!}{(d+m)!}\sum_{v\in V}B_{d+m}(t\langle v,z\rangle,{\mathbf{a}}_{v})$ (30)
+valid for almost all $z\in\mathbb{C}^{d}$, and all $t\in\mathbb{Z}_{>0}$. $\square$
+Specializing Theorem 3 to smooth polytopes immediately gives the following set of vanishing identities.
+Corollary 7.
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope.
+Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$, and for all $t\in\mathbb{Z}_{>0}$:
+  1. (a)
+    For each $0\leq m\leq d-1$, we have:
+    $0$ $=\sum_{v\in V}B_{m}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big).$ (31)
+  2. (b)
+    Equivalently, we may also express the latter set of identities by using the Barnes numbers, and removing the variable $t$.
+For each $(m,r)$ with $0\leq m\leq d-1,\text{ and }0\leq r\leq m$, and for almost all $z\in\mathbb{C}^{d}$ we have:
+    $0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}B_{m-r}({\mathbf{a}}_{v}).$ (32)
+    $\square$
+Specializing Corollary 2 to smooth polytopes, we have the following.
+Corollary 8.
+Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope. The we have:
+$1=\frac{(1)^{d}}{d!}\sum_{v\in V}B_{d}({\mathbf{a}}_{v}),$ (33)
+for almost all $z\in\mathbb{C}^{d}$. $\square$
+For example, it follows from Corollary 8 that for integer smooth polygons, we have the amusing identity:
 $12=\sum_{v\in\rm{V}}\left(\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+3\right),$ (34)
 valid for almost all $z\in\mathbb{C}^{2}$ (see Example 5).
+The integer $12$ is easy to explain - it arises naturally from the second classical Bernoulli number $B_{2}=\frac{1}{6}$.
+Similarly, we easily get a corresponding identity for all $3$-dimensional smooth polytopes:
+$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+3\right),$
+valid for almost all $z\in\mathbb{C}^{3}$ (see Example 6).
 Such identities appear to be new.
-The integer $12$ is easy to explain - it arises naturally from the second classical Bernoulli number $B_{2}=\frac{1}{6}$.
-Similarly, we have a corresponding identity for all $3$-dimensional unimodular polytopes:
-$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{1}{2}\right).$
-There is now a vast literature on Ehrhart polynomials and quasi-polynomials, and the reader may consult, for example, the books [4] [9] [21] [25] [37] [38], and papers [2] [5] [10] [12] [14] [15] [16] [17] [18] [19] [20] [22] [23] [24] [26] [29] [30] [31] [33] [34] [35] [39].
-We remark that given a triangulation of any rational polytope into partially-open simplices, one can obtain similar formulas as in this paper.
+There is now a vast literature on Ehrhart polynomials and quasi-polynomials, and the reader may consult, for example, the books [5] [10] [25] [29] [42] [43], and papers [2] [3] [6] [11] [13] [15] [16] [17] [18] [19] [20] [21] [22] [23] [24] [26] [27] [28] [30] [33] [34] [35] [37] [38] [39] [40] [44].
+We remark that given a triangulation of any rational polytope into partially-open simplices, one can obtain similar formulas to those we obtained in Theorem 1 (and consequently the other theorems here), for its Ehrhart quasi-polynomial and its discrete moments.
 However, for general rational polytopes, finding an explicit triangulation is computationally expensive, even for general $3$-dimensional polytopes: some of its vertices might have $2^{100}$ edges, for example.
-2 Known properties of Barnes polynomials
-Given their defining generating function (1), it is natural that the Barnes polynomials should inherit some of the properties of the classical Bernoulli polynomials.
-First, the Barnes polynomials can be easily given in terms of a finite sum of classical Bernoulli polynomials:
-$\frac{1}{k!}B_{k}(t,{\mathbf{a}})=\sum_{m_{1}+\cdots+m_{d}=k}\frac{a_{1}^{m_{1}-1}}{m_{1}!}\cdots\frac{a_{d}^{m_{d}-1}}{m_{d}!}B_{m_{1}}\Big(\tfrac{t}{a_{1}+\cdots+a_{d}}\Big)\cdots B_{m_{d}}\Big(\tfrac{t}{a_{1}+\cdots+a_{d}}\Big),$ (33)
-where the $B_{m}(x)$ are the classical Bernoulli polynonomials, ${\mathbf{a}}:=(a_{1},\dots,a_{d})\in\mathbb{C}^{d}$, and the indices are nonnegative integers $m_{j}\geq 0$ (see Bayad and Kim [7, Theorem 2]).
-We see that if all of the $a_{j}$’s are rational numbers, then $B_{k}(t,{\mathbf{a}})$ is a polynomial in $t$, with rational coefficients.
-From (9) and (10), it follows at once that $(a_{1}a_{2}\cdots a_{d})B_{k}(t,{\mathbf{a}})\in\mathbb{Q}[a_{1},\dots,a_{d}]$.
-Also, it’s clear that the coefficients of $B_{k}(t,{\mathbf{a}})$, as a polynomial in $t$, are symmetric functions of $a_{1},\dots,a_{d}$.
-The Barnes polynomials enjoy the following properties.
-[reciprocity law]
-$(-1)^{k}B_{k}(-t,{\mathbf{a}})=B_{k}(t+a_{1}+\cdots+a_{d},{\mathbf{a}}),$ (34)
-where ${\mathbf{a}}:=(a_{1},\dots,a_{d})^{T}$.
-[homogeneity]
-$B_{k}(\lambda t,\lambda{\mathbf{a}})=\lambda^{k-d}B_{k}(t,{\mathbf{a}}),$ (35)
-for each $k\geq 1$, and for any fixed nonzero $\lambda\in\mathbb{C}$.
-Here by definition $\lambda{\mathbf{a}}:=(\lambda a_{1},\dots,\lambda a_{d})$.
-See [7] for proofs and more background on Barnes polynomials.
-Using identity (9), we list the first few Barnes numbers with ${\mathbf{a}}\in\mathbb{C}^{d}$:
-$B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}.$ (36)
-$B_{1}({\mathbf{a}})=-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (37)
-$(a_{1}\cdots a_{d})B_{2}({\mathbf{a}})=\frac{1}{6}\sum_{k=1}^{d}a_{k}^{2}+\frac{1}{2}\sum_{1\leq i<j\leq d}a_{i}a_{j}.$ (38)
-$(a_{1}\cdots a_{d})B_{3}({\mathbf{a}})=-\frac{1}{4}\sum_{1\leq i\not=j\leq d}a_{i}^{2}a_{j}-\frac{1}{8}\sum_{1\leq i<j<k\leq d}a_{i}a_{j}a_{k}.$ (39)
-In addition, using the identities (10) and (9), we list the first few Barnes polynomials with ${\mathbf{a}}\in\mathbb{C}^{d}$:
-$B_{0}(t,{\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}},$ (40)
-$B_{1}(t,{\mathbf{a}})$ $=\binom{1}{0}B_{0}({\mathbf{a}})t+\binom{1}{1}B_{1}({\mathbf{a}})$ (41)
-$=\frac{1}{a_{1}\cdots a_{d}}t-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (42)
-$B_{2}(t,{\mathbf{a}})=\binom{2}{0}B_{0}({\mathbf{a}})t^{2}+\binom{2}{1}B_{1}({\mathbf{a}})t+\binom{2}{2}B_{2}({\mathbf{a}})$ (43)
-$=\frac{1}{a_{1}\cdots a_{d}}t^{2}-\Big(\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}\Big)t+\Big(\frac{1}{6}\sum_{k=1}^{d}\frac{a_{k}^{2}}{a_{1}\cdots a_{d}}+\frac{1}{2}\sum_{1\leq i<j\leq d}\frac{a_{i}a_{j}}{a_{1}\cdots a_{d}}\Big).$ (44)
-Finally, for the purposes of explicitly computing some Ehrhart polynomials, we list the first few relevant Barnes polynomials, for each dimension $d=1,2,3$, derived using the relations (9), (10), (36), (37), (38), and (39):
-$B_{1}(t,(a_{1}))=\frac{1}{a_{1}}t-\frac{1}{2}.$ (45)
-$B_{2}\left(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right)\right)=t^{2}\left(\frac{1}{a_{1}a_{2}}\right)-t\left(\frac{1}{a_{2}}+\frac{1}{a_{1}}\right)+\frac{1}{6}\left(\frac{a_{2}}{a_{1}}+\frac{a_{1}}{a_{2}}\right)+\frac{1}{2}.$ (46)
-$B_{3}\left(t,\left(\begin{smallmatrix}a_{1}\\ \\ a_{2}\\ \\ a_{3}\end{smallmatrix}\right)\right)=\binom{3}{0}B_{0}({\mathbf{a}})t^{3}+\binom{3}{1}B_{1}({\mathbf{a}})t^{2}+\binom{3}{2}B_{2}({\mathbf{a}})t+\binom{3}{3}B_{3}({\mathbf{a}})$ (47)
-$=\frac{1}{a_{1}a_{2}a_{3}}t^{3}-\frac{3}{2}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}}\right)t^{2}$ (48)
-$+\frac{3}{a_{1}a_{2}a_{3}}\left(\frac{1}{6}(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})+\frac{1}{2}(a_{1}a_{2}+a_{1}a_{3}+a_{2}a_{3})\right)t$ (49)
-$+\frac{1}{a_{1}a_{2}a_{3}}\Big(-\frac{1}{4}\sum_{1\leq i\not=j\leq 3}a_{i}^{2}a_{j}-\frac{1}{8}\sum_{1\leq i<j<k\leq 3}a_{i}a_{j}a_{k}\Big)$ (50)
-$=\frac{1}{a_{1}a_{2}a_{3}}t^{3}-\frac{3}{2}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}}\right)t^{2}$ (51)
-$+\frac{1}{2}\left(\left(\frac{a_{1}}{a_{2}a_{3}}+\frac{a_{2}}{a_{1}a_{3}}+\frac{a_{3}}{a_{1}a_{2}}\right)+3\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\right)t$ (52)
-$-\frac{1}{4}\left(\frac{a_{1}}{a_{2}}+\frac{a_{1}}{a_{3}}+\frac{a_{2}}{a_{1}}+\frac{a_{2}}{a_{3}}+\frac{a_{3}}{a_{1}}+\frac{a_{3}}{a_{2}}\right)-\frac{1}{8}.$ (53)
-3 Dimension 1: intervals
+2 Small dimensions
+Here we show how to specialize the formula of Theorem 1 to obtain the classical results in dimensions $1$ and $2$.
+2.1 Dimension 1: intervals
 It is instructive to consider the obligatory $1$-dimensional examples.
@@
 and here the vector of linear forms is just ${\mathbf{a}}_{v}=z$ for the vertex $v:=a$, and ${\mathbf{a}}_{v}=-z$ for the vertex $v:=b$.
-In dimension $1$ all polytopes are unimodular, so by Corollary 2 we have:
+In dimension $1$ all polytopes are smooth, so by Corollary 5 we have:
 $L_{\mathcal{P}}(t)$ $=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)=-1\cdot B_{1}(taz,z)-1\cdot B_{1}(tbz,-z)$
@@
 $\left|t{\mathcal{P}}\cap\mathbb{Z}^{d}\right|$ $=\frac{(-1)^{d}}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k}$
-$=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)-\sum_{v\in\rm{V}}B_{0}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}$
+$=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)-\sum_{v\in\rm{V}}B_{0}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\mathbb{Z}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}$
 $=-\Big(\frac{t\alpha z}{z}-\frac{1}{2}+\frac{t\beta z}{-z}-\frac{1}{2}\Big)-\Big(\frac{1}{z}\{-t\alpha\}z+\frac{1}{-z}(\{-t\beta\}-1)z\Big)$
@@
 recovering this well-known, easy result. $\square$
-4 Dimension 2: polygons
+2.2 Dimension 2: polygons
 Suppose we are given any integer polygon $\mathcal{P}$}, defined as the convex hull of its vertices $\{v_{1},\cdots v_{N}\}$.
-At each vertex $v$, we are given its edge vectors $w_{1}(v),w_{2}(v)$.
-In the ensuing formula, we define $c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v))$, and the indicator function $1_{\mathbb{Z}^{2}}(x):=\begin{cases}1&\text{if $x\in\mathbb{Z}^{2}$}\\ 0&\text{otherwise}.\end{cases}$
+At each vertex $v$, we are given its primitive edge vectors $w_{1}(v),w_{2}(v)$.
 An explicit description for the Ehrhart polynomial of $\mathcal{P}$}, using Corollary 1, is as follows:
-$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+\left(\frac{1}{2}\sum_{v\in V}\langle v,z\rangle\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)\right)t+1,$ (55)
+$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}-\frac{1}{2}\sum_{v\in V}\left(\frac{\langle v,z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle v,z\rangle}{\langle w_{2}(v),z\rangle}\right)t+1,$ (36)
 valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$.
-To show how this follows from Theorem 1, we first recall the definition ${\mathbf{a}}_{v}:=\left(\begin{smallmatrix}\langle w_{1}(v),z\rangle\\ \langle w_{2}(v),z\rangle\end{smallmatrix}\right)$ at each vertex $v$ of $\mathcal{P}$}.
-From Theorem 1, we have:
+To show how this follows from Corollary 1, we first recall the definition ${\mathbf{a}}_{v}:=\left(\begin{smallmatrix}\langle w_{1}(v),z\rangle\\ \langle w_{2}(v),z\rangle\end{smallmatrix}\right)$ at each vertex $v$ of $\mathcal{P}$}.
+By Corollary 1, we have:
 $L_{\mathcal{P}}(t)$ $=\frac{1}{2}\sum_{k=0}^{2}\binom{2}{k}\sum_{v\in V}B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}^{2-k}$
@@
 $\left.\quad\quad+\frac{1}{6}\operatorname{vol}\Pi_{v}\left(\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}\right)+\frac{1}{2}\operatorname{vol}\Pi_{v}\right).$
-To summarize, we have:
+To simplify notation, we define
+$c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v)),$
+and we summarize:
 $L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+t\sum_{v\in V}\frac{\langle v,z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\left({\displaystyle\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}{\langle p,z\rangle}}-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle\right)+1,$ (37)
@@
 We used Corollary 2 for the constant term in (37), and equation (99) for the area of the polygon.
-Recalling that $c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v))$, the center of mass of the closure of the parallelogram $\Pi_{v}$, it’s an easy exercise that summing all the points in a fundamental parallelepiped gives us:
-$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}p=\big(\operatorname{vol}\Pi_{v}+1-1_{\mathbb{Z}^{2}}(c_{v})\big)\ c_{v}.$ (57)
-(hint: translate $\Pi_{v}\cap\mathbb{Z}^{2}$ to the origin by moving its center of mass, and notice that the new non-zero points cancel in paris). So we have:
-$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}\langle p,z\rangle-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle=(1-1_{\mathbb{Z}^{2}}(c_{v}))\langle c_{v},z\rangle.$
+It’s an easy exercise that summing all the points in a $2$-dimensional fundamental parallelepiped gives us:
+$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}p=\big(\operatorname{vol}\Pi_{v}-1\big)\ c_{v}.$ (38)
+[hint: translate $\Pi_{v}\cap\mathbb{Z}^{2}$ to the origin by (its center of mass) $c_{v}$, and notice that the new non-zero points cancel in pairs]. So we have:
+$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}\langle p,z\rangle-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle=-\langle c_{v},z\rangle.$
 Substituting the latter expression into (37), we obtain
-$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+t\sum_{v\in V}\frac{\langle v,z\rangle\langle c_{v},z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)+1$
-$={\rm area}({\mathcal{P}})t^{2}+\left(\frac{1}{2}\sum_{v\in V}\langle v,z\rangle\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)\right)t+1,$
-as claimed.
+$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}-t\sum_{v\in V}\frac{\langle v,z\rangle\langle c_{v},z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}+1$ (39)
+$={\rm area}({\mathcal{P}})t^{2}-\frac{1}{2}\sum_{v\in V}\left(\frac{\langle v,z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle v,z\rangle}{\langle w_{2}(v),z\rangle}\right)t+1,$ (40)
+using the definition $\langle c_{v},z\rangle=\frac{1}{2}\left(\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle\right)$.
+Finally, let’s see how to translate the coordinate-free formulation (40) into the standard known result on Ehrhart polynomials.
+For any $x\in\mathbb{R}^{d}$, we define $\gcd(x):=|\gcd(x_{1},\dots,x_{d})|$.
+We consider any four consecutive vertices of $\mathcal{P}$}, say $v_{1},v_{2},v_{3},v_{4}$.
+To give the primitive edge vectors for $v_{2}$, we define $g_{21}:=\gcd(v_{1}-v_{2})$, which is the number of integer points on the half-open edge joining $v_{1}$ and $v_{2}$.
+We note also that, for example, $g_{23}=g_{32}$.
+Then $w_{1}(v_{2})=\frac{1}{g_{21}}(v_{1}-v_{2})$, and similarly $w_{2}(v_{2})=\frac{1}{g_{23}}(v_{3}-v_{2})$.
+In the summation of (40), we focus on just the two vertices $v_{2}$ and $v_{3}$, and we consider their contribution to the sum:
+$\left(\frac{\langle v_{2},z\rangle}{\langle w_{1}(v_{2}),z\rangle}+\frac{\langle v_{2},z\rangle}{\langle w_{2}(v_{2}),z\rangle}\right)+\left(\frac{\langle v_{3},z\rangle}{\langle w_{1}(v_{3}),z\rangle}+\frac{\langle v_{3},z\rangle}{\langle w_{2}(v_{3}),z\rangle}\right)$ (41)
+$=\frac{g_{21}\langle v_{2},z\rangle}{\langle v_{1}-v_{2},z\rangle}+\left(\frac{g_{23}\langle v_{2},z\rangle}{\langle v_{3}-v_{2},z\rangle}+\frac{g_{32}\langle v_{3},z\rangle}{\langle v_{2}-v_{3},z\rangle}\right)+\frac{g_{34}\langle v_{3},z\rangle}{\langle v_{4}-v_{3},z\rangle}$ (42)
+$=\frac{g_{21}\langle v_{2},z\rangle}{\langle v_{1}-v_{2},z\rangle}-g_{23}+\frac{g_{34}\langle v_{3},z\rangle}{\langle v_{4}-v_{3},z\rangle}.$ (43)
+Once we consider the full summation in (40), we therefore see that the simplification in pairs gives us the classical result:
+$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+\frac{1}{2}t\sum_{j=1}^{N}\gcd\left(v_{j}-v_{j+1}\right)+1,$ (44)
+where $N$ is the number of vertices of $\mathcal{P}$}, and by definition $v_{N+1}:=v_{1}$.
 Example 3.
@@
 The reason for these equal length geodesics is easy, and arises from the finite abelian group $\mathbb{Z}^{d}/\Lambda$ structure, where $\Lambda$ is the integer sublattice generated by the vertices of $\Pi_{v}$.
-Thus, for example in Figure 2, the total length of the red geodesics equals the total length of the blue geodesics, because the blue geodesics arise from an index $2$ subgroup of $\mathbb{Z}^{2}/\Lambda$ here.
+Thus, for example in Figure 2, there are exactly two closed geodesics, and the total length of the red geodesic equals the total length of the blue geodesic, because here the blue geodesic arises from an index $2$ subgroup of $\mathbb{Z}^{2}/\Lambda$.
 $\square$
 Example 4.
-In $\mathbb{R}^{2}$, just for fun let’s retrive the Ehrhart polynomial for the standard triangle ${\mathcal{P}}:=\operatorname*{conv}\{\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\}$, using Barnes polynomials.
-We’ll use Corollary 2 of Theorem 1, since the standard triangle is a unimodular polygon.
-Recalling that the only Barnes polynomials that occur in dimension $2$ are:
+In $\mathbb{R}^{2}$, just for fun let’s retrieve the Ehrhart polynomial for the standard triangle ${\mathcal{P}}:=\operatorname*{conv}\{\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\}$, using Barnes polynomials.
+We’ll use Corollary 5 of Theorem 1, since the standard triangle is a smooth polygon.
+We recall that the only Barnes polynomials that occur in dimension $2$ are:
 $B_{0}(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right))=\frac{1}{a_{1}a_{2}},\quad\quad B_{1}(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right))=\frac{1}{a_{1}a_{2}}t-\frac{1}{2}\frac{a_{1}+a_{2}}{a_{1}a_{2}},$ (45)
@@
 $v_{3}:=\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\implies\omega_{1}(v_{3})=\left(\begin{smallmatrix}\ 1\\ -1\end{smallmatrix}\right),\omega_{2}(v_{3})=\left(\begin{smallmatrix}\ 0\\ -1\end{smallmatrix}\right).$
-Hence by definition, we have ${\mathbf{a}}_{v_{1}}:=\Big(\langle z,\omega_{1}(v_{1})\rangle,\langle z,\omega_{2}(v_{1})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}\\ z_{2}\end{smallmatrix}\right)$.
-Similarly, by definition we have ${\mathbf{a}}_{v_{2}}:=\Big(\langle z,\omega_{1}(v_{2})\rangle,\langle z,\omega_{2}(v_{2})\rangle\Big)^{T}=\left(\begin{smallmatrix}-z_{1}+z_{2}\\ -z_{1}\end{smallmatrix}\right)$, and
-${\mathbf{a}}_{v_{3}}:=\Big(\langle z,\omega_{1}(v_{3})\rangle,\langle z,\omega_{2}(v_{3})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}-z_{2}\\ -z_{2}\end{smallmatrix}\right)$.
+Therefore by definition, we have ${\mathbf{a}}_{v_{1}}:=\Big(\langle z,\omega_{1}(v_{1})\rangle,\langle z,\omega_{2}(v_{1})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}\\ z_{2}\end{smallmatrix}\right)$.
+Similarly ${\mathbf{a}}_{v_{2}}:=\Big(\langle z,\omega_{1}(v_{2})\rangle,\langle z,\omega_{2}(v_{2})\rangle\Big)^{T}=\left(\begin{smallmatrix}-z_{1}+z_{2}\\ -z_{1}\end{smallmatrix}\right)$, and ${\mathbf{a}}_{v_{3}}:=\Big(\langle z,\omega_{1}(v_{3})\rangle,\langle z,\omega_{2}(v_{3})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}-z_{2}\\ -z_{2}\end{smallmatrix}\right)$.
 By Corollary 5 , we have
@@
 and we see that all of the $z$’s miraculously cancel out, as predicted by Theorem 1. $\square$
-5 The unit cube in dimension d
-We consider the $d$-dimensional unit cube $\square:=[-1,1]^{d}$, another unimodular polytope.
+2.3 A d-dimensional cube
+We consider the $d$-dimensional cube $\square:=[-1,1]^{d}$, another smooth polytope.
 Although it is trivial to compute the Ehrhart polynomial in this case, namely $L_{\square}(t)=(2t+1)^{d}$, the verification will shed some light on our use of the Barnes polynomials, and even gives new identities (equation (48) below) for the discrete hypercube $\{-1,1\}^{d}$.
@@
 for each $0\leq j\leq d$, and for almost all $z\in\mathbb{C}^{d}$.
-We may notice that the case $j=0$ of identity (61) was already covered by Corollary 5, in the special case of unimodular polytopes.
+We may notice that the case $j=0$ of identity (48) was already covered by Corollary 2, in the special case of smooth polytopes.
 The case $j=d$ of identity (48) seems interesting, and using $B_{0}({\mathbf{a}}_{v})=\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}$, tells us that
@@
 $\sum_{n\in\mathcal{K}_{v}\cap\mathbb{Z}^{d}}e^{\langle n,z\rangle}=\frac{\sigma_{\Pi_{v}+v}(z)}{\prod_{k=1}^{d}\left(1-e^{\langle\omega_{k},z\rangle}\right)}.$ (54)
-[9, Theorem 3.4].
+[10, Theorem 3.4].
 We use the notation $\mathcal{K}_{v}-v:=\mathcal{K}_{0}(v)$, the translation of the cone $\mathcal{K}_{v}$ to the origin.
@@
 In other words, we used the definition of the Barnes polynomial, but with the vector ${\mathbf{a}}_{v}:=\Big(\langle w_{1}(v),z_{0}\rangle,\dots,\langle w_{d}(v),z_{0}\rangle\Big)^{T}$.
-To recap, using (69), we may sum the latter contributions (73), one from each of the vertex tangent cones, to obtain:
+To summarize, by using (56) we may sum the latter contributions (60), one from each of the vertex tangent cones, to obtain:
 $\sigma_{t{\mathcal{P}}}(xz_{0})$ $=\sigma_{t\mathcal{K}_{v_{1}}}(xz_{0})+\cdots+\sigma_{t\mathcal{K}_{v_{N}}}(xz_{0})$
@@
 Namely, we use $B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j}$, in equation (74): (68):
-$\frac{1}{m!}\mu_{m}(t{\mathcal{P}},z_{0})=(-1)^{d}\sum_{v\in V}\sum_{k=0}^{d+m}\sum_{j=0}^{k}\tfrac{1}{k!(d+m-k)!}\binom{k}{j}B_{j}({\mathbf{a}}_{v})t^{k-j}\langle v,z_{0}\rangle^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-k},$ (88)
+$\tfrac{(-1)^{d}}{m!}\mu_{m}(t{\mathcal{P}},z_{0})=\sum_{v\in V}\sum_{k=0}^{d+m}\sum_{j=0}^{k}\tfrac{1}{k!(d+m-k)!}\binom{k}{j}B_{j}({\mathbf{a}}_{v})t^{k-j}\langle v,z_{0}\rangle^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-k},$ (75)
 Finally, we just have to sum over all relevant $k,j$, such that kj:=rk-j:=r, with $r$ fixed, so that we may isolate the coefficient of $t^{r}$ on the right-hand side.
@@
 In our context, (75) therefore implies that the coefficient of $t^{r}$ equals:
-$(-1)^{d}\sum_{v\in V}\sum_{j=0}^{d+m-r}\frac{1}{(d+m-r-j)!j!r!}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j}$ (89)
-$=\frac{(-1)^{d}}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j},$ (90)
-valid for almost all $z_{0}\in\mathbb{C}^{d}$.
-To prove part (c), we simply differentiate the latter expression in $t$:
-$\frac{d}{dt}d_{r+1}(t)=\frac{d}{dt}\left(\tfrac{(-1)^{d}m!}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-r-j}\right)$
-$=(-1)^{d}m!\sum_{v\in V}\langle v,z\rangle^{r+1}\sum_{j=0}^{d-r-1}\tfrac{(d+m-r-j)}{r!j!(d+m-r-j)!}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d+m-r-j-1}$
+$\sum_{v\in V}\sum_{j=0}^{d+m-r}\frac{1}{(d+m-r-j)!j!r!}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j}$ (76)
+$=\frac{1}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{d+m-r-j},$ (77)
+valid for almost all $z_{0}\in\mathbb{C}^{d}$. To prove part (c), we simply differentiate the latter expression in $t$:
+$\frac{d}{dt}d_{r}(t)=\frac{d}{dt}\left(\tfrac{m!}{(d+m)!}\sum_{v\in V}\sum_{j=0}^{d+m-r}\binom{d+m}{r,j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{r}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{d+m-r-j}\right)$
+$=m!\sum_{v\in V}\langle v,z\rangle^{r+1}\sum_{j=0}^{d-r-1}\tfrac{(d+m-r-j)}{r!j!(d+m-r-j)!}B_{j}({\mathbf{a}}_{v})\sum_{q\in(\Pi_{v}+tv)\cap\mathbb{Z}^{d}}{\Big(\langle q,z\rangle-t\langle v,z\rangle\Big)}^{d+m-r-j-1}$
 $=-(r+1)d_{r+1}(t),$
-8 Proof of Theorem 3: vanishing identities for integer simple polytopes
+5 Proof of Theorem 3: vanishing identities for rational simple polytopes
 We may push the proof of Theorem 1 a bit further, to derive certain canonical vanishing identities for simple integer polytopes $\mathcal{P}$}, arising from the singular part of (66), which is a Laurent series in $x$.
@@
 (of Theorem 3)Using equation (66) from the proof of Theorem 1, we recall:
-$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in\rm{V}}\sigma_{\Pi_{v}}(xz_{0})\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}.$
+$\sigma_{t{\mathcal{P}}}(xz_{0})$ $=(-1)^{d}\sum_{v\in V}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}\sum_{n,k\geq 0}{\langle q,z_{0}\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}$ (78)
 To prove part (a), we begin by noting that the left-hand-side of the latter identity is an entire function of $x$.
 We therefore know that for each fixed $n<0$, the coefficient of any $x^{n}$ on the right-hand-side must be identically zero.
-We next expand the exponential function $e^{x\langle z_{0},p\rangle}$ as a Taylor series as well, obtaining:
-$(-1)^{d}\sigma_{t{\mathcal{P}}}(xz_{0})=$ $\sum_{v\in V}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}e^{x\langle z_{0},p\rangle}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (91)
-$=\sum_{v\in V}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\sum_{n\geq 0}{\langle z_{0},p\rangle}^{n}\frac{x^{n}}{n!}\sum_{k\geq 0}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k-d}}{k!}$ (92)
-$=\sum_{v\in V}\sum_{n,k\geq 0}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle z_{0},p\rangle}^{n}B_{k}(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v})\frac{x^{k+n-d}}{k!n!}$ (93)
-The left-hand side of (91) is an entire function of $x$.
-Therefore we cannot have any singularities in $x$, so that comparing the coefficient of $x^{m-d}$ on both sides, for each fixed value of $m\in\{0,\dots,d-1\}$, we must get vanishing identities.
+We now compare the coefficient of $x^{m-d}$ on both sides, and for each fixed value of $m\in\{0,\dots,d-1\}$, we must get vanishing identities.
 Precisely, we get the following $d$ identities for each integer polytope in $\mathbb{R}^{d}$, by summing over all $k,n\geq 0$ in (78) with $k+n=m$ fixed:
-$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z_{0}\rangle}^{m-k}.$ (94)
+$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z_{0}\rangle,{\mathbf{a}}_{v}\Big)\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-k}.$ (79)
 To prove part (b), we simplify (79) further by noticing that we have an identity for all positive integer values of $t$.
@@
 We may use the explicit polynomial representation of the Barnes-Bernoulli polynomials (10), namely $B_{k}(t,{\mathbf{a}})=\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}})t^{k-j}$, together with (79), implies that
-$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in\rm{V}}\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{k-j}t^{k-j}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}$ (95)
-$=\sum_{v\in\rm{V}}\sum_{k=0}^{m}\sum_{j=0}^{k}\binom{m}{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{k-j}t^{k-j}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}$ (96)
+$0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in\rm{V}}\sum_{j=0}^{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{k-j}t^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-k}$ (80)
+$=\sum_{v\in\rm{V}}\sum_{k=0}^{m}\sum_{j=0}^{k}\binom{m}{k}\binom{k}{j}B_{j}({\mathbf{a}}_{v})\langle v,z_{0}\rangle^{k-j}t^{k-j}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-k}$ (81)
 Now we just have to fix kj:=rk-j:=r, and set the coefficient of $t^{r}$ equal to zero. We may use the following elementary combinatorial identity, when setting kj=rk-j=r:
@@
 In our context, (81) therefore implies, after a little simplification:
-$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\frac{{\langle p,z\rangle}^{m-r-j}}{(m-r-j)!},$ (97)
+$0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!(m-r-j)!}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-r-j}$ (82)
 for each $(m,r)$ with $0\leq m\leq d-1,\text{ and }0\leq r\leq m$. ∎
-9 Proof of Corollary 2: Ehrhart for unimodular polytopes
+6 Proof of Corollary 5: Ehrhart for smooth polytopes
 Proof.
-(of Corollary 2)The unimodularity assumption means that for all vertices $v\in{\mathcal{P}}$, there is exactly one integer point in the fundamental parallelepiped $\Pi_{v}$, namely the origin.
+(of Corollary 5)The smooth polytope assumption means that for all vertices $v\in{\mathcal{P}}$, there is exactly one integer point in the fundamental parallelepiped $\Pi_{v}$, namely the origin.
 Therefore the discrete moments, for each $0\leq k\leq d-1$, degenerate to:
@@
 Example 5.
-In dimension $d=2$, Corollary 5 had an immediate consequence, namely (31), which tells us that for any unimodular integer polygon $P\subset\mathbb{R}^{2}$ we have:
+In dimension $d=2$, Corollary 8 tells us that for any smooth integer polygon $P\subset\mathbb{R}^{2}$ we have:
 $1=\frac{1}{2}\sum_{v\in\rm{V}}B_{2}({\mathbf{a}}_{v})=\frac{1}{2}\sum_{v\in\rm{V}}\left(\frac{1}{6}\left(\frac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}\right)+\frac{1}{2}\right),$ (84)
@@
 Example 6.
-In dimension $d=3$, the same consequence (31) of Corollary 5 tells us that for any unimodular integer polytope $P\subset\mathbb{R}^{3}$, we have:
-$1=\frac{(-1)^{3}}{3!}\sum_{v\in\rm{V}}B_{3}({\mathbf{a}}_{v}),$
-and unpacking $B_{3}({\mathbf{a}}_{v})$ by using (53), we arrive at the identity
-$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{1}{2}\right),$
+In dimension $d=3$, Corollary 8 tells us that for any smooth integer polytope $P\subset\mathbb{R}^{3}$, we have:
+$1=\frac{(-1)^{3}}{3!}\sum_{v\in\rm{V}}B_{3}({\mathbf{a}}_{v})=-\frac{1}{6}\sum_{v\in\rm{V}}\frac{1}{8a_{1}(v)a_{2}(v)a_{3}(v)}(s_{1}s_{2}-s_{1}^{3}),$
+by using (113) of Appendix A. Using the definitions of the power sums $s_{j}$ and defining $a_{j}(v):=\langle w_{j}(v),z\rangle$, we arrive at the identity
+$24$ $=\frac{1}{2}\sum_{v\in\rm{V}}\frac{s_{1}^{3}-s_{1}s_{2}}{a_{1}(v)a_{2}(v)a_{3}(v)}$ (86)
+$=\frac{1}{2}\sum_{v\in\rm{V}}\tfrac{\left(\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle+\langle w_{3}(v),z\rangle\right)^{3}-\left(\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle+\langle w_{3}(v),z\rangle\right)\left(\langle w_{1}(v),z\rangle^{2}+\langle w_{2}(v),z\rangle^{2}+\langle w_{3}(v),z\rangle^{2}\right)}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle\langle w_{3}(v),z\rangle}$ (87)
+$=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+3\right),$ (88)
 valid for almost all $z\in\mathbb{C}^{3}$. $\square$
 7.1 Special cases of Theorem 3
-We expand on a few of the vanishing identities from Theorem 3.
+Restricting attention here to integer polytopes, we expand on a few of the vanishing identities from Theorem 3.
 We first note that the particular cases $r=m$ recovers known identities, by Brion and Vergne [13] and that the other identities appear to be new.
@@
 Here we used $B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}$, and $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}1=|\Pi_{v}\cap\mathbb{Z}^{d}|=\operatorname{vol}\Pi_{v}$.
-The latter identity is a standard fact, which may be arrived using Fourier transforms of polytopes [12] [37, Theorem 5.11].
+The latter identity is a standard fact from the geometry of numbers [42, Theorem 5.11].
+Throughout this section, we fix the dimension $d\geq 2$.
 Regarding the new identities, we first discuss the case $m=1,r=0$, as it presents a special case that is of independent interest: equation (94) below.
@@
 $\sum_{v\in\rm{V}}\frac{1}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=\frac{1}{2}\sum_{v\in\rm{V}}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}.$ (94)
-Case (m,r)=(m,m1)(m,r)=(m,m-1), with $2\leq m\leq d$. Here Theorem 3 gives us
+Case (m,r)=(m,m1)(m,r)=(m,m-1), with $1\leq m\leq d-1$. Here Theorem 3, for the special case of integer simple polytopes, gives us
 $0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\sum_{j=0}^{1}\frac{B_{j}({\mathbf{a}}_{v})}{j!}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\frac{{\langle p,z\rangle}^{1-j}}{(1-j)!}$ (95)
@@
 $\sum_{v\in\rm{V}}\frac{\langle v,z\rangle^{m-1}}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=\frac{1}{2}\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\operatorname{vol}\Pi_{v}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle},$ (97)
-for all simple integer polytopes, for almost all $z\in\mathbb{C}^{d}$, and for each $1\leq m\leq d$.
-Next, we make a simple observation that holds for the family of unimodular polytopes.
-If $\mathcal{P}$} is a unimodular polytope, then the latter identity (109) enjoys a simplification.
-For a unimodular polytope, we have $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=0$, because the origin is the only integer point in $\Pi_{v}$.
+for all simple integer polytopes, for almost all $z\in\mathbb{C}^{d}$.
+Next, we make a simple observation that holds for the family of smooth polytopes.
+If $\mathcal{P}$} is a smooth polytope, then the latter identity (97) enjoys a simplification.
+For a smooth polytope, we have $\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\langle p,z\rangle=0$, because the origin is the only integer point in $\Pi_{v}$.
 Therefore
 $\sum_{v\in\rm{V}}\langle v,z\rangle^{m-1}\frac{\sum_{k=1}^{d}\langle w_{k}(v),z\rangle}{\prod_{k=1}^{d}\langle w_{k}(v),z\rangle}=0,$ (98)
-for each fixed $1\leq m\leq d$, and for almost all $z\in\mathbb{C}^{d}$.
+for each fixed $1\leq m\leq d-1$, and for almost all $z\in\mathbb{C}^{d}$.
 Remark 8.
@@
 $\operatorname{vol}{\mathcal{P}}=\frac{(-1)^{d}}{d!}\sum_{v\in\rm{V}}\left(\operatorname{vol}\Pi_{v}\right)\langle v,z\rangle^{d}B_{0}({\mathbf{a}}_{v})=\frac{(-1)^{d}}{d!}\sum_{v\in\rm{V}}\left(\operatorname{vol}\Pi_{v}\right)\frac{\langle v,z\rangle^{d}}{\prod_{m=1}^{d}\langle w_{m}(v),z\rangle},$ (99)
-for almost all $z\in\mathbb{C}^{2}$. We’ve recovered here a known volume identity of Brion [11, Corollary 2] and Lawrence [27] (see also [37, Theorem 7.23] for another proof).
+for almost all $z\in\mathbb{C}^{2}$. We’ve recovered here a known volume identity of Brion [12, Corollary 2] and Lawrence [31] (see also [42, Theorem 7.23] for a Fourier-style proof).
 8 Further remarks and research directions
@@
 For simplicity, in this section we write the discrete moments for $t=1$ as $\mu_{n}(\Pi_{v})$.
-Conjecture 1.
-Fix the dimension $d$, and a half-open integer paralleleppiped $\Pi\subset\mathbb{R}^{d}$.
-Then for each fixed integer $k$, and fixed $z\in\mathbb{C}^{d}$, there is a polynomial-time algorithm, in the bit-input of the vertices of an integer parallelepiped $\Pi$, for computing the discrete moment $\mu_{k}(\Pi)$.
+Question 1.
+Fix the dimension $d$, and fix a half-open integer paralleleppiped $\Pi\subset\mathbb{R}^{d}$, defined by its edges $w_{1},\dots,w_{d}$.
+Fix also a positive integer $k$ and $z\in\mathbb{C}^{d}$.
+Let $N$ be the bit input of the edges $w_{1},\dots,w_{d}$.
+Is there an $N^{O(d)}$-time algorithm for computing the discrete moment $\mu_{k}(\Pi)$?
+Barvinok’s algorithm [6] guarantees a complexity bound of $N^{O(d\log d)}$ for general rational polytopes, where $N$ is the bit input of $\mathcal{P}$}.
+So in particular Barvinok’s algorithm guarantees a polynomial-time algorithm for computing $\mu_{k}(\Pi)$ in fixed dimension.
+But with half-open integer parallelepipeds we have a lot more structure, suggesting a slightly better bound.
 Remark 9.
-We may give a trivial bound for the discrete moments $\mu_{n}(\Pi_{v})$ in Theorem 1, and hence for the coefficients of Ehrhart quasi-polynomials by considering $M_{v}:=\max_{p\in\Pi_{v}}\|p\|$.
+We may give a trivial bound for the discrete moments $\mu_{n}(\Pi_{v})$ in Theorem 1, and hence for the coefficients of Ehrhart quasi-polynomials by considering $M_{v}:=\displaystyle\max_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}\|p\|$.
 We have
@@
 Of course, the latter bound is far from optimal.
-Question 1.
+Question 2.
 What is a good bound for the discrete moments $\mu_{n}(\Pi_{v})$?
@@
 Let $d\geq 2$.
-We know from Section 2 that $B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v})\prod_{j=1}^{d}\langle w_{j},z\rangle$ is a nonconstant polynomial in $z_{1},\dots,z_{d}$, for all $k\geq 1$.
+We know from §A that $\left(\prod_{j=1}^{d}\langle w_{j},z\rangle\right)B_{k}\left((t\langle v,z\rangle,{\mathbf{a}}_{v}\right)$ is a nonconstant polynomial in $z_{1},\dots,z_{d}$, for all $k\geq 1$.
 Picking a generic $z\in\mathbb{C}^{d}$, in the sense of (7), already implies that $z\in\mathbb{C}^{d}$ satisfies $\prod_{j=1}^{d}\langle w_{j},z\rangle\not=0$.
 Now pick a generic $z\in\mathbb{C}^{d}$ that also belong to the variety defined by
-$\bigcap_{k=1}^{d-1}\left\{z\in\mathbb{C}^{d}\ \mid\ B_{k}(t\langle v,z\rangle,{\mathbf{a}}_{v}).\prod_{j=1}^{d}\langle w_{j},z\rangle=0\right\}.$ (112)
-This variety gives us extra flexibility to choose points $z\in\mathbb{C}^{d}$ that allow us to omit the discrete moment sums $\mu_{k}(\Pi_{v})$ (for $k=1,2,\dots,d-1$) from our calculations.
-Due to the flexibility in picking any generic $z\in\mathbb{C}^{d}$ in the theorems, we arrive at the following conclusion:
+$X:=\bigcap_{k=1}^{d-1}\left\{z\in\mathbb{C}^{d}\ \mid\ \left(\prod_{j=1}^{d}\langle w_{j},z\rangle\right)B_{k}\left((t\langle v,z\rangle,{\mathbf{a}}_{v}\right)=0\right\}.$ (100)
+The variety $X$ in (100), as a vanishing set of polynomials in $z\in\mathbb{C}^{d}$, gives us extra flexibility to choose points $z\in\mathbb{C}^{d}$ that allow us to omit the discrete moment sums $\mu_{k}(\Pi_{v})$ (for $k=1,2,\dots,d-1$) from our calculations.
+Due to the flexibility in picking any generic $z\in\mathbb{C}^{d}$ in the theorems, we arrive at the following conclusion, if we know that the variety (100) is nonempty:
 In computing the Ehrhart coefficients $c_{1},\dots c_{d-1}$ of Corollary 1, we may ignore any one fixed vertex of our choice.
-Question 2.
-Can we ignore two or more vertices, in computing the Ehrhart coefficients $c_{1},\dots c_{d-1}$ of Corollary 1, by using the variety in Remark 10?
+Question 3.
+Can we ignore two or more vertices, in computing the Ehrhart coefficients $c_{1},\dots c_{d-1}$ of Corollary 1, by using the variety $X$ of Remark 10?
+In small dimensions, it is easy to show that the variety $X$ is nonempty. In general, we inquire about its nonempty behavior.
+Question 4.
+Is the variety $X$ in (100) nonempty for all dimensions $d$?
+Question 5.
+Can we understand complete period collapse for $d$-dimensional rational polytopes by understanding the lattice flows of the fundamental parallelepipeds at the vertices, appearing in Theorem 1 ?
+Appendix A Known properties of Barnes polynomials
+Given their defining generating function (1), it is natural that the Barnes polynomials should inherit some of the properties of the classical Bernoulli polynomials.
+First, the Barnes polynomials can be easily given in terms of a finite sum of classical Bernoulli polynomials:
+$\frac{1}{k!}B_{k}(t,{\mathbf{a}})=\sum_{m_{1}+\cdots+m_{d}=k}\frac{a_{1}^{m_{1}-1}}{m_{1}!}\cdots\frac{a_{d}^{m_{d}-1}}{m_{d}!}B_{m_{1}}\Big(\tfrac{t}{a_{1}+\cdots+a_{d}}\Big)\cdots B_{m_{d}}\Big(\tfrac{t}{a_{1}+\cdots+a_{d}}\Big),$ (101)
+where the $B_{m}(x)$ are the classical Bernoulli polynonomials, ${\mathbf{a}}:=(a_{1},\dots,a_{d})\in\mathbb{C}^{d}$, and the indices are nonnegative integers $m_{j}\geq 0$ (see Bayad and Kim [8, Theorem 2]).
+We see that if all of the $a_{j}$’s are rational numbers, then $B_{k}(t,{\mathbf{a}})$ is a polynomial in $t$, with rational coefficients.
+From (9) and (10), it follows at once that $(a_{1}a_{2}\cdots a_{d})B_{k}(t,{\mathbf{a}})\in\mathbb{Q}[a_{1},\dots,a_{d}]$.
+Also, it’s clear that the coefficients of $B_{k}(t,{\mathbf{a}})$, as a polynomial in $t$, are symmetric functions of $a_{1},\dots,a_{d}$.
+The Barnes polynomials enjoy the following properties.
+  1. (a)
+    [Reciprocity]
+    $(-1)^{k}B_{k}(-t,{\mathbf{a}})$ $=B_{k}(t+a_{1}+\cdots+a_{d},{\mathbf{a}}).$
+  2. (b)
+    [Homogeneity]
+    $B_{k}(\lambda t,\lambda{\mathbf{a}})=\lambda^{k-d}B_{k}(t,{\mathbf{a}}).$
+  3. (c)
+    [Differentiation]
+    $\frac{d}{dt}B_{k}(t;\mathbf{a})=kB_{k-1}(t;\mathbf{a}).$
+The properties above hold for each $k\geq 1$, and for any fixed nonzero $\lambda\in\mathbb{C}$, where by definition $\lambda{\mathbf{a}}:=(\lambda a_{1},\dots,\lambda a_{d})$, and ${\mathbf{a}}:=(a_{1},\dots,a_{d})^{T}$.
+See [8] for proofs and more background on Barnes polynomials.
+For the purposes of explicitly computing some Ehrhart polynomials, it is very useful to list the first few Barnes polynomials in low dimensions, relevant for us due to the theorems above.
+The first five Barnes polynomials in dimension d:
+$B_{0}(t,{\mathbf{a}})$ $=\frac{1}{a_{1}\cdots a_{d}}.$ (102)
+$B_{1}(t,{\mathbf{a}})$ $=\frac{1}{a_{1}\cdots a_{d}}t-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (103)
+$B_{2}(t,{\mathbf{a}})$ $=\frac{1}{a_{1}\cdots a_{d}}\left(t^{2}-t\sum_{i=1}^{d}a_{i}-\frac{1}{12}\sum_{i=1}^{d}a_{i}^{2}+\frac{1}{4}\left(\sum_{i=1}^{d}a_{i}\right)^{2}\right).$ (104)
+Alternatively, the standard power sum notation helps us write the formulas more compactly, especially as the dimension increases: $s_{r}:=\displaystyle\sum_{j=1}^{d}a_{j}^{r}$.
+$(a_{1}\cdots a_{d})B_{2}(t,{\mathbf{a}})$ $=t^{2}-s_{1}t+\frac{1}{12}\left(3s_{1}^{2}-s_{2}\right).$ (105)
+$(a_{1}\cdots a_{d})B_{3}(t,{\mathbf{a}})$ $=t^{3}-\frac{3}{2}s_{1}t^{2}+\frac{1}{4}\left(3s_{1}^{2}-s_{2}\right)t+\frac{1}{8}\left(s_{1}s_{2}-s_{1}^{3}\right)$ (106)
+$=\left(t-\frac{1}{2}s_{1}\right)^{3}-\frac{1}{4}s_{2}\left(t-\frac{1}{2}s_{1}\right).$ (107)
+$(a_{1}\cdots a_{d})B_{4}(t,{\mathbf{a}})$ $=t^{4}-2s_{1}t^{3}+\left(\tfrac{3s_{1}^{2}-s_{2}}{2}\right)t^{2}-\left(\tfrac{s_{1}^{3}-s_{1}s_{2}}{2}\right)t+\tfrac{15s_{1}^{4}-30s_{1}^{2}s_{2}+10s_{2}^{2}+14s_{4}}{240}$ (108)
+$=\left(t-\frac{1}{2}s_{1}\right)^{4}-\frac{1}{2}s_{2}\left(t-\frac{1}{2}s_{1}\right)^{2}+\frac{1}{24}s_{2}^{2}+\frac{7}{240}s_{4}.$ (109)
+Remark 11.
+We remark that it’s straightforward to prove the following property of Barnes polynomials, via a symmetry argument in their generating function.
+For odd $k$ the Barnes polynomials $B_{k}(t,{\mathbf{a}})$ are odd polynomials in the variable $\left(t-\frac{1}{2}s_{1}\right)$, while for even $k$ the Barnes polynomials $B_{k}(t,{\mathbf{a}})$ are even polynomials in the variable $\left(t-\frac{1}{2}s_{1}\right)$.
+The first five Barnes numbers in dimension $d$:
+$B_{0}({\mathbf{a}})$ $=\frac{1}{a_{1}\cdots a_{d}}.$ (110)
+$B_{1}({\mathbf{a}})$ $=-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (111)
+$(a_{1}\cdots a_{d})B_{2}({\mathbf{a}})$ $=\frac{1}{12}\left(3s_{1}^{2}-s_{2}\right).$ (112)
+$(a_{1}\cdots a_{d})B_{3}({\mathbf{a}})$ $=\frac{1}{8}(s_{1}s_{2}-s_{1}^{3}).$ (113)
+$(a_{1}\cdots a_{d})B_{4}({\mathbf{a}})$ $=\frac{1}{16}s_{1}^{4}-\frac{1}{8}s_{1}^{2}s_{2}+\frac{1}{24}s_{2}^{2}+\frac{7}{120}s_{4}.$ (114)
+For the purposes of computing examples of Ehrhart quasi-polynomials in $\mathbb{R}$ and $\mathbb{R}^{2}$, we record the explicit special cases of $d=1$ and $d=2$ here.
+Barnes polynomials in dimension 1, relevant for us:
+$B_{0}(t,(a_{1}))=\frac{1}{a_{1}},\quad B_{1}(t,(a_{1}))=\frac{1}{a_{1}}t-\frac{1}{2}.$ (115)
+Barnes polynomials in dimension 2, relevant for us:
+$B_{0}(t,{\mathbf{a}})$ $=\frac{1}{a_{1}a_{2}}$ (116)
+$B_{1}(t,{\mathbf{a}})$ $=\frac{1}{a_{1}a_{2}}t-\frac{1}{2}\frac{a_{1}+a_{2}}{a_{1}a_{2}}.$ (117)
+$B_{2}\left(t,{\mathbf{a}}\right)$ $=t^{2}\left(\frac{1}{a_{1}a_{2}}\right)-t\left(\frac{1}{a_{2}}+\frac{1}{a_{1}}\right)+\frac{1}{6}\left(\frac{a_{2}}{a_{1}}+\frac{a_{1}}{a_{2}}\right)+\frac{1}{2}.$ (118)
+Barnes numbers in dimension 2, relevant for us:
+$B_{0}({\mathbf{a}})=\frac{1}{a_{1}a_{2}},\quad B_{1}({\mathbf{a}})=-\frac{1}{2}\frac{a_{1}+a_{2}}{a_{1}a_{2}},\quad B_{2}({\mathbf{a}})=\frac{1}{6}\left(\frac{a_{2}}{a_{1}}+\frac{a_{1}}{a_{2}}\right)+\frac{1}{2}.$ (119)